### v1

# Advancing Local Search in SMT-NRA with MCSAT Integration

Abstract.
In this paper, we advance local search for Satisfiability Modulo the Theory of Nonlinear Real Arithmetic (SMT-NRA for short). First, we introduce a two-dimensional cell-jump move, called $2d$-cell-jump, generalizing the key operation, cell-jump, of the local search method for SMT-NRA. Then, we propose an extended local search framework, named $2d$-LS (following the local search framework, LS, for SMT-NRA), integrating the model constructing satisfiability calculus (MCSAT) framework to improve search efficiency. To further improve the efficiency of MCSAT, we implement a recently proposed technique called sample-cell projection operator for MCSAT, which is well suited for CDCL-style search in the real domain and helps guide the search away from conflicting states. Finally, we design a hybrid framework for SMT-NRA combining MCSAT, $2d$-LS and OpenCAD, to improve search efficiency through information exchange. The experimental results demonstrate improvements in local search performance, highlighting the effectiveness of the proposed methods.
SMT-NRA, Local Search, MCSAT, Hybrid Method
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Theory of computation Automated reasoning
1. Introduction
Satisfiability Modulo Theories (SMT) is concerned with determining the satisfiability of first-order logic formulas under background theories, such as integer arithmetic, real arithmetic, arrays, bit vectors, strings, and others. This paper concentrates on SMT problems over the theory of quantifier-free nonlinear real arithmetic (NRA), referred to as SMT-NRA. The goal is to determine the satisfiability of polynomial formulas, which are expressed in the form of $\bigwedge_{i}\bigvee_{j}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0,\text{where}~{}\rhd_{ij}\in\{<,>,\leq,\geq,=,\neq\}$ and $p_{ij}(\bar{\boldsymbol{x}})$ are polynomials. SMT-NRA has found widespread applications in various fields, including for example control theory for system verification (Alur, 2011; Cimatti et al., 2013b; Chen and He, 2018), robotics for motion planning and trajectory optimization (Nedunuri et al., 2014; Shoukry et al., 2016; Imeson and Smith, 2019), and software/hardware verification to ensure timing and performance constraints in embedded systems (Trindade and Cordeiro, 2016; Beyer et al., 2018; Li and Gopalakrishnan, 2010; Faure-Gignoux et al., 2024). It also plays a critical role in optimization (Sebastiani and Tomasi, 2012; Li et al., 2014; Bjørner et al., 2015), where nonlinear constraints are common.
Tarski proposed an algorithm in 1951 (Tarski, 1998), solving the problem of quantifier elimination (QE) of the first-order theory over real closed fields, which takes SMT-NRA as a special case. Cylindrical Algebraic Decomposition (CAD), another real QE method introduced by Collins in 1975 (Collins, 1975), can solve polynomial constraints by decomposing space into finitely many regions (called cells) arranged cylindrically. CAD provides a more practical approach to quantifier elimination than Tarski’s procedure though it remains of doubly exponential complexity. In practice, the Model-Constructing Satisfiability Calculus (MCSAT) (de Moura and Jovanović, 2013) is a widely used complete SMT algorithms. MCSAT integrates two solvers from the classical framework into one solver that simultaneously searches for models in both the Boolean structure and the theory structure, thereby constructing consistent Boolean assignments and theory assignments. Several state-of-the-art (SOTA) SMT solvers supporting NRA have been developed over the past two decades. Representative SOTA solvers include Z3 (de Moura and Bjørner, 2008) and Yices2 (Dutertre, 2014), which implement MCSAT, as well as CVC5 (Barbosa et al., 2022) and MathSAT5 (Cimatti et al., 2013a), which use alternative techniques.
The computational complexity of SMT-NRA remains a challenge, motivating research of incomplete solvers that are usually more efficient in finding SAT assignments. Local search, a popular paradigm, has been developed in recent years for SMT-NRA (Li et al., 2023; Li and Cai, 2023; Wang et al., 2024). Local search begins with a theory assignment and approaches a model of the polynomial formula iteratively by moving locally. This process ends when a model is found or other termination conditions are met. The most effective move for real-space search is the ‘cell-jump’ proposed by Li et al. (Li et al., 2023), which leads sample points to different CAD cells via one-dimensional moves.
The complementarity between complete methods and local search has led to the development of hybrid solvers for related problems. Notable examples include the hybridization of Conflict-Driven Clause Learning (CDCL) and local search for SAT (Letombe and Marques-Silva, 2008; Cai and Zhang, 2021), as well as the combination of CDCL(T) and local search (Zhang et al., 2024) for solving satisfiability modulo the theory of nonlinear integer arithmetic, SMT-NIA for short. These studies suggest that a similar hybrid approach may hold promise for solving SMT-NRA.
In this paper, we aim at advancing local search in SMT-NRA on problems in the form of
$\bigwedge_{i}\bigvee_{j}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0,\text{where}~{}\rhd_{ij}\in\{<,>,\neq\}~{}\text{and}~{}p_{ij}\in\mathbb{Q}[\bar{\boldsymbol{x}}],$
through the integration of MCSAT and make the following contributions:
  • $\bullet$
    We propose a new cell-jump mechanism, called $2d$-cell-jump, which supports two-dimensional search and may find models faster.
  • $\bullet$
    We propose an extended local search framework, named $2d$-LS, integrating the MCSAT framework to improve search efficiency.
  • $\bullet$
    Inspired by the work (Zhang et al., 2024) of Zhang et al. on the combination of CDCL(T) and local search for SMT-NIA, we design a hybrid framework for SMT-NRA that exploits the complementary strengths of MCSAT, $2d$-LS and OpenCAD (Han et al., 2014). In this framework, MCSAT drives $2d$-LS to accelerate the search for a model, $2d$-LS helps MCSAT identify unsatisfiable cells, and OpenCAD is utilized to handle unsatisfiable formulas dominated by algebraic conflicts.
  • $\bullet$
    The above proposed methods have been implemented as a solver called HELMS. When implementing MCSAT, we use a recently proposed technique called sample-cell projection operator for MCSAT, which further improves the efficiency of MCSAT. Comparison to SOTA solvers on a large number of benchmarks shows that the newly proposed methods are effective.
The rest of this paper is organized as follows. Section 2 introduces preliminaries of the problem in SMT-NRA, the local search solver and the sample-cell projection for MCSAT. Section 3 extends local search to $2d$-LS, introducing the new cell-jump. Section 4 outlines the hybrid method that combines $2d$-LS, MCSAT and OpenCAD. The experimental results in Section 5 demonstrate the progress on local search. Finally, Section 6 concludes this paper.
2. Preliminaries
2.1. Problem Statement
Let $\bar{\boldsymbol{x}}=(x_{1},\ldots,x_{n})$ be a vector of variables. Denote by $\mathbb{Q}$, $\mathbb{R}$, $\mathbb{Z}$ and $\mathbb{N}$ the set of rational numbers, real numbers, integer numbers and natural numbers, respectively. Let $\mathbb{Q}[\bar{\boldsymbol{x}}]$ be the ring of polynomials in the variables $x_{1},\ldots,x_{n}$ with coefficients in $\mathbb{Q}$.
Definition 2.1 (Polynomial Formula).
The following formula
(1) $F=\bigwedge_{i=1}^{M}\bigvee_{j=1}^{m_{i}}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$
is called a polynomial formula, where $1\leq i\leq M<+\infty,1\leq j\leq m_{i}<+\infty$, $p_{ij}\in\mathbb{Q}[\bar{\boldsymbol{x}}]$ and $\rhd_{ij}\in\{<,>,\leq,\geq,=,\neq\}$. Moreover, $\bigvee_{j=1}^{m_{i}}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$ is called a polynomial clause or simply a clause, and $p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$ is called an atomic polynomial formula or simply an atom.
For any polynomial formula $F$, a complete assignment is a mapping $\alpha:\bar{\boldsymbol{x}}\to\mathbb{R}^{n}$ such that $x_{1}\mapsto a_{1},\ldots,x_{n}\mapsto a_{n}$, where every $a_{i}\in\mathbb{R}$. We denote by $\alpha[x_{i}]$ the assigned value $a_{i}$ of the variable $x_{i}$. With slight abuse of notation, we sometimes represent a complete assignment simply by the real vector $(a_{1},\ldots,a_{n})$. An atom is true under $\alpha$ if it evaluates to true under $\alpha$, and otherwise it is false under $\alpha$, A clause is satisfied under $\alpha$ if at least one atom in the clause is true under $\alpha$, and falsified under $\alpha$ otherwise. When the context is clear, we simply say a true (or false) atom and a satisfied (or falsified) clause. A polynomial formula is satisfiable (SAT) if there exists a complete assignment in $\mathbb{R}^{n}$ such that all clauses in the formula are satisfied, and such an assignment is a model to the polynomial formula. A polynomial formula is unsatisfiable if any assignment is not a model.
Example 2.2.
(A running example) We take the following polynomial formula as a running example
$F_{r}=f_{1,r}<0\land f_{2,r}<0,$
where $r\in\mathbb{N}$, $f_{1,r}=x^{2}+y_{1}^{2}+\cdots+y_{r}^{2}-z^{2}$ and $f_{2,r}=(x-3)^{2}+y_{1}^{2}+\cdots+y_{r}^{2}+z^{2}-5$. Let $\ell_{1,r}$ denote atom $f_{1,r}<0$ and $\ell_{2,r}$ denote atom $f_{2,r}<0$. Under the assignment $(x,y_{1},\ldots,y_{r},z)\mapsto(\frac{3}{2},0,\ldots,0,\frac{8}{5})$, both $\ell_{1,r}$ and $\ell_{2,r}$ are true, and thus $F_{r}$ is satisfiable.
The problem we consider in this paper is to determine the satisfiability of polynomial formulas in the form of (1) with $\rhd_{ij}\in\{<,>,\neq\}.$
2.2. Cell-Jump Operation in Local Search
Li et al. propose a local search algorithm (Li et al., 2023, Alg. 3) for solving SMT-NRA. The key point of the algorithm is the cell-jump operation (Li et al., 2023, Def. 11 & Alg. 2), which updates the current assignment along a straight line with given direction.
  • $\bullet$
    Cell-Jump Along a Coordinate Axis Direction: The first type of cell-jump moves along one coordinate axis direction to update the current assignment. For instance, consider an SMT-NRA problem with two variables $x$ and $y$. Note that the search space is $\mathbb{R}^{2}$. This type of cell-jump moves either along the $x$-axis direction, i.e., updating the assigned value of the first variable $x$, or along the $y$-axis direction, i.e., updating the assigned value of the second variable $y$.
  • $\bullet$
    Cell-Jump Along a Given Direction: The second type of cell-jump moves along any given direction. For instance, consider an SMT-NRA problem with two variables. Given a straight line with the direction $(3,4)$, one cell-jump moves from assignment $(a_{1},a_{2})$ to a new assignment $(a_{1}+3t,a_{2}+4t)$ along the line. Such movement enables a more comprehensive exploration of the search space, potentially facilitating the rapid discovery of solutions.
2.3. Sample-Cell Projection Operation for MCSAT
In our MCSAT implementation, we use the projection operator, called sample-cell projection operator proposed in (Li and Xia, 2020), which is essentially the same as the ‘biggest cell’ heuristic in the recent work (Nalbach et al., 2024). Compared with (Brown, 2013), the sample-cell projection operator is better suited for integration with the MCSAT framework, enabling both efficient solving and lazy evaluation. This subsection briefly introduces the sample-cell projection operator.
Let $f,g\in\mathbb{Q}[\bar{\boldsymbol{x}}]$, $F$ be a finite subset of $\mathbb{Q}[\bar{\boldsymbol{x}}]$ and $a\in\mathbb{R}^{n}$. Denote by $\mathop{\mathtt{disc}}(f,x_{i})$ and $\mathop{\mathtt{res}}(f,g,x_{i})$ the discriminant of $f$ with respect to $x_{i}$ and the resultant of $f$ and $g$ with respect to $x_{i}$, respectively. The order of $f$ at $a$ is defined as
${\mathop{\mathtt{order}}}_{a}(f)=\min($$\{k\in\mathbb{N}\mid\text{some partial derivative of total order }k\text{ of }f$
[…]
We call $f$ order-invariant on $S\subseteq\mathbb{R}^{n}$, if ${\mathop{\mathtt{order}}}_{a_{1}}(f)={\mathop{\mathtt{order}}}_{a_{2}}(f)$ for any $a_{1},a_{2}\in S$. Note that order-invariance implies sign-invariance. We call $F$ a square-free basis in $\mathbb{Q}[\bar{\boldsymbol{x}}]$, if the elements in $F$ are of positive degrees, primitive, square-free and pairwise relatively prime.
Definition 2.3 (Analytic Delineable).
(Collins, 1975) Let $r\geq 1$, $S$ be a connected sub-manifold of $\mathbb{R}^{r-1}$ and $f\in\mathbb{Q}[x_{1},\ldots,x_{r}]\setminus\{0\}$. The polynomial $f$ is called analytic delineable on $S$, if there exist finitely many analytic functions $\theta_{1},\ldots,\theta_{k}:S\rightarrow\mathbb{R}$ (for $k\geq 0$) such that
  • $\bullet$
    $\theta_{1}<\cdots<\theta_{k}$,
  • $\bullet$
    the set of real roots of the univariate polynomial $f(a,x_{r})$ is $\{\theta_{1}(a),\ldots,\theta_{k}(a)\}$ for all $a\in S$, and
  • $\bullet$
    there exist positive integers $m_{1},\ldots,m_{k}$ such that for any $a\in S$ and for $j=1,\ldots,k$, the multiplicity of the root $\theta_{j}(a)$ of $f(a,x_{r})$ is $m_{j}$.
Let sample point $a=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and $F=\{f_{1},\ldots,f_{s}\}\subseteq\mathbb{Q}[\bar{\boldsymbol{x}}]\setminus\{0\}$. Consider the real roots of univariate polynomials in
(2) $\{f_{1}(a_{1},\ldots,a_{n-1},x_{n}),\ldots,f_{s}(a_{1},\ldots,a_{n-1},x_{n})\}\setminus\{0\}.$
Denote by $\gamma_{i,k}~{}(\in\mathbb{R})$ the $k$-th real root of $f_{i}(a_{1},\ldots,a_{n-1},x_{n})$. We define the sample polynomial set of $F$ at $a$, denoted by $\mathop{\mathtt{s\_poly}}(F,x_{n},a)$, as follows.
  1. (1)
    If there exists $\gamma_{i,k}$ such that $\gamma_{i,k}=a_{n}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i}\}$.
  2. (2)
    If there exist two real roots $\gamma_{i_{1},k_{1}}$ and $\gamma_{i_{2},k_{2}}$ such that $\gamma_{i_{1},k_{1}}<a_{n}<\gamma_{i_{2},k_{2}}$ and the open interval $(\gamma_{i_{1},k_{1}},\gamma_{i_{2},k_{2}})$ contains no $\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i_{1}},f_{i_{2}}\}$.
  3. (3)
    If there exists $\gamma_{i^{\prime},k^{\prime}}$ such that $a_{n}>\gamma_{i^{\prime},k^{\prime}}$ and for all $\gamma_{i,k}$, $\gamma_{i^{\prime},k^{\prime}}\geq\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i^{\prime}}\}$.
  4. (4)
    If there exists $\gamma_{i^{\prime},k^{\prime}}$ such that $a_{n}<\gamma_{i^{\prime},k^{\prime}}$ and for all $\gamma_{i,k}$, $\gamma_{i^{\prime},k^{\prime}}\leq\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i^{\prime}}\}$.
  5. (5)
    Specially, if every polynomial in (2) has no real roots, define $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\emptyset$.
For every $f\in F$, suppose $f=c_{m}x_{n}^{d_{m}}+c_{m-1}x_{n}^{d{m-1}}+\cdots+c_{0}x_{n}^{d_{0}}$, where every $c_{i}\in\mathbb{Q}[x_{1},\ldots,x_{n-1}]\setminus\{0\}$, $d_{i}\in\mathbb{N}$ and $d_{m}>d_{m-1}>\cdots>d_{0}$. If there exists $j\in\mathbb{N}$ such that $c_{j}(a_{1},\ldots,a_{n-1})\neq 0$ and $c_{j^{\prime}}(a_{1},\ldots,a_{n-1})=0$ for any $j^{\prime}>j$, then define the sample coefficients of $f$ at $a$ as $\mathop{\mathtt{s\_coeff}}(f,x_{n},a)=\{c_{m},c_{m-1},\ldots,c_{j}\}$. Otherwise, define $\mathop{\mathtt{s\_coeff}}(f,x_{n},a)=\{c_{m},c_{m-1},\ldots,c_{j}\}$
Definition 2.4 (Sample-Cell Projection).
(Li and Xia, 2020) Suppose $F$ is a finite polynomial subset of $\mathbb{Q}[\bar{\boldsymbol{x}}]\setminus\{0\}$ and $a=(a_{1},\ldots,a_{n-1})\in\mathbb{R}^{n-1}$. The sample-cell projection of $F$ on $x_{n}$ at $a$ is defined as
$\mathop{\mathtt{Proj}}(F,x_{n},a)=$ $\bigcup_{f\in F}\{\mathop{\mathtt{s\_coeff}}(f,x_{n},a)\}\cup\bigcup_{f\in F}\{\mathop{\mathtt{disc}}(f,x_{n})\}$
$\cup\bigcup_{\begin{subarray}{c}f\in F,\\ g\in\mathop{\mathtt{s\_poly}}(F,x_{n},a),\\ f\neq g\end{subarray}}\{\mathop{\mathtt{res}}(f,g,x_{n})\}.$
Theorem 2.5.
(Li and Xia, 2020) Let $n\geq 2$, $F$ be a square-free basis in $\mathbb{Q}[\bar{\boldsymbol{x}}]$, $a=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$, and $S$ be a connected sub-manifold of $\mathbb{R}^{n-1}$ such that $(a_{1},\ldots,a_{n-1})\in S$. If every element of $\mathop{\mathtt{Proj}}(F,x_{n},a)$ is order-invariant on $S$, then every element of $F$ either vanishes identically on $S$ or is analytic delineable on $S$.
3. Extending Local Search with MCSAT
In Section 3.1, we propose a new cell-jump operation, called $2d$-cell-jump. Comparing to cell-jump in (Li et al., 2023, Alg. 2), $2d$-cell-jump allows searching for a model in a plane instead of along a line. In Section 3.2, based on $2d$-cell-jump, we develop a new local search algorithm for SMT-NRA, called $2d$-LS. The algorithm can be considered as an extension of LS (Li et al., 2023, Alg. 3).
3.1. New Cell-Jump: $2d$-Cell-Jump
Remark that the cell-jump operation proposed in (Li et al., 2023) is limited to searching for a solution along a straight line, which is one-dimensional. With the assistance of MCSAT, the search process can be extended into higher dimensional space. In this subsection, we define $2$-dimensional cell-jump, $2d$-cell-jump for short, expanding the cell-jump move from a line parallel to a coordinate axis to a plane parallel to an axes plane, and from a given line to a given plane. Theoretically, this approach can be generalized to $D$-dimensional space, where $D\geq 2$. For the sake of MCSAT’s efficiency, we adopt $D=2$ in this paper.
3.1.1. New Sample Point
Note that sample points (Li et al., 2023, Def. 10) are candidate assignments to move to in original cell-jump operation. To define the new cell-jump operation, we first introduce new sample points. These sample points are generated using MCSAT, where we seek one model (if exists) for atomic polynomial formulas by fixing all variables except for two specific ones.
Definition 3.1 (Sample Point).
Let $n\geq 2$. Consider atom $\ell:p(\bar{\boldsymbol{x}})>0$ (or atom $\ell:p(\bar{\boldsymbol{x}})<0$), and suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. For any pair of distinct variables $x_{i}$ and $x_{j}$ $(i<j)$, let $p^{*}(x_{i},x_{j})=p(a_{1},\ldots,a_{i-1},x_{i},a_{i+1},\ldots,a_{j-1},x_{j},a_{j+1},\ldots,a_{n})\allowbreak\in\mathbb{Q}[x_{i},x_{j}]$. If $p^{*}(x_{i},x_{j})>0$ (or $p^{*}(x_{i},x_{j})<0$) is satisfiable and $(b_{i},b_{j})\in\mathbb{Q}^{2}$ is a model of it, then $(a_{1},\ldots,a_{i-1},b_{i},a_{i+1},\ldots,a_{j-1},b_{j},a_{j+1},\ldots,a_{n})$ is a sample point of $\ell$ with respect to (w.r.t.) $x_{i},x_{j}$ under $\alpha$.
Remark that a sample point of an atom is a model of it. In practice, we use MCSAT to determine the satisfiability of $p^{*}(x_{i},x_{j})>0$ (or $p^{*}(x_{i},x_{j})<0$). The reason is that MCSAT can quickly determine the satisfiability of two-variable polynomial formulas. If a formula only contains one variable, the formula can be solved directly by real root isolation.
Example 3.2.
Consider atoms $\ell_{1,r}:f_{1,r}<0$ and $\ell_{2,r}:f_{2,r}<0$ in Example 2.2. Suppose the current assignment is $\alpha:(x,y_{1},\ldots,y_{r},z)\mapsto(0,0,\ldots,0,0)$. Keeping the variables $x$ and $z$, we have $f_{1,r}|\allowbreak_{y_{1}=0,\ldots,y_{r}=0}=x^{2}-z^{2}$, and $(1,2)$ is a model of it. So, $(1,0,\ldots,0,2)$ is a sample point of $\ell_{1,r}$ w.r.t. $x,z$ under $\alpha$. Keeping the variables $y_{r}$ and $z$, we obtain $f_{2,r}|\allowbreak_{x=0,y_{1}=0,\ldots,y_{r-1}=0}=y_{r}^{2}+z^{2}+4$, which has no models. So, there is no sample point of $\ell_{2,r}$ w.r.t. $y_{r},z$ under $\alpha$.
Let $r=1$. We have $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}$, $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5$ and $\alpha:(x,y_{1},z)\mapsto(0,0,0)$. As shown in Fig. 1, the region above the orange cone satisfies $\ell_{1,1}$, and that inside the blue sphere satisfies $\ell_{2,1}$. So, every point in the intersection of the region above the orange cone and the green plane is a sample point of $\ell_{1,1}$ w.r.t. $x,z$ under $\alpha$, such as point $(1,0,2)$. The region inside the blue sphere and the red plane has no intersection. So, there is no sample point of $\ell_{2,1}$ w.r.t. $y_{r},z$ under $\alpha$.
Figure 1. The figure of a sample point/$2d$-cell-jump operation in a plane parallel to an axes plane. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes $\{(x,0,z)\mid x\in\mathbb{R},z\in\mathbb{R}\}$, and the red plane denotes $\{(0,y_{1},z)\mid y_{1}\in\mathbb{R},z\in\mathbb{R}\}$. The current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$ (i.e., the vertex of the yellow cone). Point $(1,0,2)$ is a sample point of $\ell_{1,1}$ w.r.t. $x,z$ under $\alpha$. There is a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{1},e_{3})$ operation jumping from $(0,0,0)$ to $(1,0,2)$ in the green plane.
3.1.2. New Cell-Jump
For any $i$ $(1\leq i\leq n)$, let $e_{i}=(0,\ldots,1,\allowbreak\ldots,0)$ be a vector in $\mathbb{R}^{n}$ with $1$ in the $i$-th position. For any point $\alpha=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and any two linearly independent vectors $d_{1},d_{2}\in\mathbb{R}^{n}$, let
$\alpha+\langle d_{1},d_{2}\rangle=\{a_{1}e_{1}+\cdots+a_{n}e_{n}+\lambda_{1}d_{1}+\lambda_{2}d_{2}\mid\lambda_{1}\in\mathbb{R},\lambda_{2}\in\mathbb{R}\},$
which is a plane spanned by vectors $d_{1}$ and $d_{2}$, passing through point $\alpha$. Specially, for any $i,j$ $(1\leq i<j\leq n)$, $(0,\ldots,0)+\langle e_{i},e_{j}\rangle$ is called an axes plane. And $\alpha+\langle e_{i},e_{j}\rangle$ is a plane parallel to an axes plane.
Definition 3.3 ($2d$-Cell-Jump in a Plane Parallel to an Axes Plane).
Suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. Let $\ell$ be a false atom $p(\bar{\boldsymbol{x}})<0$ or $p(\bar{\boldsymbol{x}})>0$. For each pair of distinct variables $x_{i}$ and $x_{j}$ $(i<j)$ such that there exists a sample point $\alpha_{s}$ of $\ell$ w.r.t. $x_{i},x_{j}$ under $\alpha$, there exists a $2d$-cell-jump operation in the plane $\alpha+\langle e_{i},e_{j}\rangle$, denoted as ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell,\alpha,e_{i},e_{j})$, updating $\alpha$ to $\alpha_{s}$ (making $\ell$ become true).
Example 3.4.
Consider polynomial formula $F_{r}=f_{1,r}<0\land f_{2,r}<0$ in Example 2.2. Suppose the current assignment is $\alpha:(x,y_{1},\ldots,y_{r},z)\mapsto(0,0,\ldots,0,0)$. Both atoms $\ell_{1,r}:f_{1,r}<0$ and $\ell_{2,r}:f_{2,r}<0$ are false under $\alpha$. Recalling Example 3.2, $(1,0,\ldots,0,2)$ is a sample point of $\ell_{1,r}$ w.r.t. $x,z$ under $\alpha$, and there is no sample point of $\ell_{2,r}$ w.r.t. $y_{r},z$ under $\alpha$. So, there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,r},\alpha,e_{1},e_{r+2})$ operation, updating $\alpha$ to $(1,0,\ldots,0,2)$, and no ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,r},\alpha,e_{r+1},e_{r+2})$ operation exists. Let $r=1$. As shown in Fig. 1, the ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{1},e_{3})$ operation jumps from $(0,0,0)$ to $(1,0,2)$ in the green plane, and there is no ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,e_{2},e_{3})$ operation in the red plane.
Example 3.5.
(a) one cell-jump move
(b) one $2d$-cell-jump move
(c) two cell-jump moves
(d) two $2d$-cell-jump moves
Figure 2. Compare cell-jump (Li et al., 2023, Def. 11) and $2d$-cell-jump. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes $\{(x,0,z)\mid x\in\mathbb{R},z\in\mathbb{R}\}$, and the red plane denotes $\{(0,y_{1},z)\mid y_{1}\in\mathbb{R},z\in\mathbb{R}\}$. The vertex of the yellow cone $(0,0,0)$ is the current assignment.
Let $r=1$ and consider polynomial formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$ in Example 2.2, where $\ell_{1,1}$ denotes atom $f_{1,1}<0$ and $\ell_{2,1}$ denotes atom $f_{2,1}<0$. Suppose the current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$. In Fig. 2, we compare the cell-jump operation $\mathop{\mathtt{cjump}}$ along a line parallel to a coordinate axis (Li et al., 2023, Def. 11) and the $2d$-cell-jump operation ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}$ in a plane parallel to an axes plane.
The vertex of the yellow cone $(0,0,0)$ is the current assignment. From point $(0,0,0)$, there exists a $\mathop{\mathtt{cjump}}(z,\ell_{1,1})$ operation jumping to $(0,0,2)$ along the $z$-axis (as shown in Fig. 2), and there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{2},e_{3})$ jumping to the same point in the $y_{1}z$-plane (as shown in Fig. 2). After a one-step move, both cell-jump and $2d$-cell-jump make $\ell_{1,1}$ become true.
Note that $(0,0,2)$ is not a model of $\ell_{2,1}$. Consider cell-jump and $2d$-cell-jump of $\ell_{2,1}$ from point $(0,0,2)$. There exists a $\mathop{\mathtt{cjump}}(z,\ell_{2,1})$ operation jumping to $(\frac{5}{2},0,2)$ along the $x$-axis (as shown in Fig. 2), while there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,e_{1},e_{2})$ jumping to $(\frac{3}{2},0,\frac{8}{5})$ in the $xy$-plane (as shown in Fig. 2). Both the second jumps make $\ell_{2,1}$ become true. It is easy to check that $(\frac{5}{2},0,2)$ is not a model of $\ell_{1,1}$, but $(\frac{3}{2},0,\frac{8}{5})$ is. So, After a two-step move, $2d$-cell-jump finds a model of formula $F_{1}$, while cell-jump does not. The reason is that the $2d$-cell-jump operation searches in a plane, covering a wider search area, potentially leading to faster model discovery.
Definition 3.6 ($2d$-Cell-Jump in a Given Plane).
Suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. Let $\ell$ be a false atom $p(\bar{\boldsymbol{x}})<0$ (or $p(\bar{\boldsymbol{x}})>0$). Given two linearly independent vectors $d_{1}=(d_{1,1},\ldots,d_{1,n})$ and $d_{2}=(d_{2,1},\ldots,d_{2,n})$ in $\mathbb{Q}^{n}$, introduce two new variables $t_{1},t_{2}$ and replace every $x_{i}$ with $a_{i}+d_{1,i}t_{1}+d_{2,i}t_{2}$ in $p(\bar{\boldsymbol{x}})$ to obtain a bivariate polynomial $p^{*}(t_{1},t_{2})$. If there exists a sample point $(t_{1}^{*},t_{2}^{*})$ of $p^{*}(t_{1},t_{2})<0$ (or $p^{*}(t_{1},t_{2})>0$) w.r.t. $t_{1},t_{2}$ under assignment $(t_{1},t_{2})\mapsto(0,0)$, then there exists a $2d$-cell-jump operation in the plane $\alpha+\langle d_{1},d_{2}\rangle$, denoted as ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell,\alpha,d_{1},d_{2})$, updating $\alpha$ to $\alpha+t_{1}^{*}d_{1}+t_{2}^{*}d_{2}$ (making $\ell$ become true).
Example 3.7.
Figure 3. The figure of a $2d$-cell-jump operation in a given plane. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes the plane $\langle(0,1,0),(15,0,16)\rangle$. The current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$ (i.e., the vertex of the yellow cone). From $(0,0,0)$, there is a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,(0,1,0),(15,0,16))$ operation jumping to $(\frac{3}{2},0,\frac{8}{5})$ in the green plane, which is a model to formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$.
Let $r=1$ and consider polynomial formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$ in Example 2.2, where $\ell_{1,1}$ denotes atom $f_{1,1}<0$ and $\ell_{2,1}$ denotes atom $f_{2,1}<0$. Suppose the current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$, and $\ell_{2,1}$ is false under $\alpha$. Given two linearly independent vectors $d_{1}=(0,1,0)$ and $d_{2}=(15,0,16)$, consider $2d$-cell-jump operations of $\ell_{2,1}$ in the plane $\alpha+\langle d_{1},d_{2}\rangle$ (the green plane in Fig. 3). Replacing $(x,y_{1},z)$ with $(15t_{2},t_{1},16t_{2})$ in $f_{2,1}$, we get a bivariate polynomial $f_{2,1}^{*}=t_{1}^{2}+481t_{2}^{2}-90t_{2}+4$. It is easy to check that $(0,\frac{1}{10})$ is a sample point of $f_{2,1}^{*}<0$. So, there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,d_{1},d_{2})$ operation, updating $\alpha$ to $\alpha+t_{1}^{*}d_{1}+t_{2}^{*}d_{2}=(\frac{3}{2},0,\frac{8}{5})$. As shown in Fig. 3, from current location $(0,0,0)$, the ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,d_{1},d_{2})$ operation jumps to point $(\frac{3}{2},0,\frac{8}{5})$ in the green plane. In fact, $(\frac{3}{2},0,\frac{8}{5})$ is not only a model to $\ell_{2,1}$ but also a model to formula $F_{1}$.
3.2. New Local Search Algorithm: $2d$-LS
Input: $F$, a polynomial formula with variables $x_{1},\ldots,x_{n}$ such that the relational operator of every atom is ‘$<$’ or ‘$>$’; $MaxRestart$, $MaxJump$ and $m$, three positive integers
Output: $status$, ‘SAT’ or ‘UNKNOWN’; $\alpha$, an assignment; $numJump$, a positive integer
1 $numRestart,numJump\leftarrow 1$
2while $numRestart\leq MaxRestart$ do
3       $\alpha\leftarrow(a_{1},\ldots,a_{n})$, an initial complete assignment in $\mathbb{Q}^{n}$
4      while $numJump\leq MaxJump$ do
5             if $F(\alpha)=$True then
6                  return ‘SAT’, $\alpha$, $numJump$
7            if there exists a cell-jump operation (Li et al., 2023, Def. 11) along a line parallel to a coordinate axis with positive score [^1] The scoring function and weighting scheme are from (Li et al., 2023). Their role is to assess candidate sample points. A positive score means that the operation of updating an assignment moves closer to assignments that satisfy the given polynomial formula, with higher scores indicating closer.  then
8                   perform such an operation with the highest score to update $\alpha$
9
10            else
11                   generate $2m$ random vectors $d_{1},\ldots,d_{2m}$, where $d_{i}\in\mathbb{Q}^{n}$
12                  $L\leftarrow\{\alpha+\langle d_{1}\rangle,\ldots,\alpha+\langle d_{2m}\rangle\}$, herein $\alpha+\langle d_{i}\rangle=\{a_{1}e_{1}+\cdots+a_{n}e_{n}+\lambda d_{i}\mid\lambda\in\mathbb{R}\}$ is a random line passing through $\alpha$ with direction $d_{i}$
13                  if there exists a cell-jump operation (Li et al., 2023, Alg. 2) along a line in $L$ with positive score then
14                         perform such an operation with the highest score to update $\alpha$
15
16                  else if there exists a $2d$-cell-jump operation in a plane parallel to an axes plane with positive score then
17                        perform such an operation with the highest score to update $\alpha$
18                   else
19                         $P\leftarrow\{\alpha+\langle d_{1},d_{2}\rangle,\ldots,\alpha+\langle d_{2m-1},d_{2m}\rangle\}$, where every $\alpha+\langle d_{i},d_{j}\rangle$ is a random plane
20                        if there exists a $2d$-cell-jump operation in a plane in $P$ with positive score then
21                              perform such an operation with the highest score to update $\alpha$
22
23                        else
24                              break
25
26
27
28            $numJump$++
29      $numRestart$++
return ‘UNKNOWN’, $\alpha$, $numJump$
Algorithm 1 $2d$-LS
Based on the new cell-jump operation, we develop a new local search algorithm, named $2d$-LS. In fact, $2d$-LS is a generalization of LS (Li et al., 2023, Alg. 3).
Recall that LS adopts a two-step search framework. Building upon LS, $2d$-LS utilizes a four-step search framework, where the first two steps are the same as LS. These four steps are described as follows, also see Algorithm 1 for reference.
  1. (1)
    Try to find a cell-jump operation (Li et al., 2023, Def. 11) along a line that passes through the current assignment point with a coordinate axis direction.
  2. (2)
    If the first step fails, generate $2m$ random vectors $d_{1},\ldots,d_{2m}$, where $m\geq 1$. Attempt to perform a cell-jump operation (Li et al., 2023, Alg. 2) along a random line, which passes through the current assignment point with direction $d_{i}$.
  3. (3)
    If the second step fails, try to find a $2d$-cell-jump operation in a plane that passes through the current assignment point and is parallel to an axes plane.
  4. (4)
    If the third step fails, attempt to perform a $2d$-cell-jump operation in a random plane, which is spanned by vectors $d_{i},d_{j}$ and passes through the current assignment point.
Note that cell-jump/$2d$-cell-jump operations are performed on false atoms. False atoms can be found in both falsified clauses and satisfied clauses, where the former are of greater significance in satisfying a polynomial formula. Thus, in every step above, we adopt the two-level heuristic in (Cai et al., 2022, Section 4.2). First, try to perform a cell-jump/$2d$-cell-jump operation to make false atoms in falsified clauses become true. If such operation does not exist, then seek one to correct false atoms in satisfiable clauses. (Due to space constraints, this two-level heuristic is not explicitly written in Algorithm 1.)
Moreover, there are two noteworthy points in $2d$-LS. First, the algorithm employs the restart mechanism to avoid searching nearby one initial assignment point. The maximum restart times $MaxRestart$ is set in the outer loop. Second, in every restart, we limit the times of cell-jumps to avoid situations where there could be infinite cell-jumps between two sample points. The maximum cell-jump times $MaxJump$ is set in the inner loop.
4. A Hybrid Framework for SMT-NRA
Inspired by the work of Zhang et al. (Zhang et al., 2024) on combining CDCL(T) and local search for SMT-NIA, we design a hybrid framework for SMT-NRA that exploits the complementary strengths of $2d$-LS (see Alg. 1), MCSAT (de Moura and Jovanović, 2013) and OpenCAD (Han et al., 2014).
The hybrid framework (also see Alg. 2) consists of the following three main stages.
  • $\bullet$
    Stage 1: $2d$-LS. Given a polynomial formula such that every relational operator appearing in it is ‘$<$’ or ‘$>$’, the first stage (Alg. 2, line 2–line 2) tries to solve the satisfiability by calling the $2d$-LS algorithm, that is Alg. 1. This stage has a 1 second time limit. The reason for employing the local search algorithm first lies in its lightweight and incomplete property. If it successfully finds a model, the solving time is typically short. If it fails, the subsequent stages apply complete solving algorithms. In addition, two key outputs from $2d$-LS are maintained for later stages: the final cell-jump location and the number of cell-jump steps (Alg. 2, line 2). The final cell-jump location provides candidate variable assignments for the second stage, while the number of cell-jump steps helps to estimate the number of unsatisfiable cells for the input formula, which will be used in the third stage.
  • $\bullet$
    Stage 2: $2d$-LS-Driven MCSAT. The second stage (Alg. 2, line 2–line 2) adopts an MCSAT framework as its foundational architecture. Recall that original MCSAT framework (de Moura and Jovanović, 2013) assigns variables one-by-one without imposing constraints on variable assignments. For example, for the theory of nonlinear real arithmetic, every variable can be assigned an arbitrary rational number. However, in Alg. 2, following each variable assignment in the MCSAT framework, we add a heuristic condition (Alg. 2, line 2) to determine whether the input formula may be satisfiable under the current variable assignments. If the formula is determined to have a high probability of being satisfiable, we invoke the $2d$-LS algorithm (i.e., Alg. 1) for the formula with all variables assigned so far substituted by their assigned values in MCSAT. The invoked $2d$-LS has a 1 second time limit.
    1. (1)
      If the output is ‘SAT’, combining current variable assignments in MCSAT and the model found by Alg. 1, we obtain a model for the original input formula.
    2. (2)
      Otherwise, we use the final cell-jump location of Alg. 1 as candidate variable assignments for the unassigned variables in MCSAT, provided it keeps consistent.
    This stage employs a local search algorithm to efficiently solving sub-formula satisfiability under partial assignments and drive variable assignments in the MCSAT framework, achieving an organic integration of the two methods. Hence, we call the stage “$2d$-LS-driven MCSAT”. For further implementation details of our MCSAT framework, please refer to Section 4.1. Additionally, to prepare for the third stage, we collect learned clauses generated during the MCSAT procedure (see Alg. 2, line 2 and line 2), and updates the estimation for the number of unsatisfiable cells (see Alg. 2, line 2).
  • $\bullet$
    Stage 3: OpenCAD. Recall that in the first two stages, we estimate the number of unsatisfiable cells for the input formula (note this represents a rough approximation, rather than the exact number of unsatisfiable cells). During the second stage, once the estimation exceeds a predetermined threshold (see Alg. 2, line 2), the algorithm transitions to this stage by invoking the OpenCAD procedure. OpenCAD is a complete method for deciding satisfiability of quantifier-free formulas comprising exclusively strict inequality constraints. The integration of OpenCAD into our hybrid framework is motivated by the complementary strengths of the two complete methods: MCSAT is a CDCL-style search framework, demonstrating superior performance on determining the satisfiability of satisfiable formulas or unsatisfiable ones dominated by Boolean conflicts, while OpenCAD specializes in efficiently handling unsatisfiable formulas dominated by algebraic conflicts (refer to Example 4.1 for a more detailed explanation). Moreover, we provide learned clauses from the MCSAT procedure for the OpenCAD invocation. The lifting process of OpenCAD terminates immediately upon detecting any low-dimensional region that violates either the input formula or these learned clauses.
Example 4.1.
Consider the following polynomial formula
$F=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2})^{2}-4(x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2}+x_{3}^{2}x_{4}^{2}+x_{4}^{2}x_{5}^{2}+x_{5}^{2}x_{1}^{2})>0.$
The Boolean structure of this formula is remarkably simple, with a single polynomial constraint. However, the algebraic structure of the only polynomial in the formula is highly complex. Given the polynomial in $F$ as input, a CAD algorithm partitions $\mathbb{R}^{5}$ into approximately $2000$ cells, on which the polynomial has constant sign, either $+$, -- or 00. In the MCSAT framework, it is tedious to encode every unsatisfiable cell of $F$. Thus, for satisfiability solving of formula $F$, OpenCAD demonstrates superior efficiency compared to MCSAT.
4.1. MCSAT Implementation
Note that Algorithm 2 consists of two parts: the black lines represent our implementation of the MCSAT framework (de Moura and Jovanović, 2013), while the blue lines indicate our hybrid framework’s extensions to the original MCSAT framework. In this subsection, we provide a detailed exposition of our MCSAT implementation.
The original MCSAT framework is described by transition relations between search states as in (Jovanović and De Moura, 2013; de Moura and Jovanović, 2013). In order to present the hybrid framework more clearly, we provide a pseudocode representation of our MCSAT implementation in black lines of Algorithm 2. We first explain some notations used in Algorithm 2.
  • $\bullet$
    $CS_{i}$ refers to the set of clauses in $F$ on level $i$, i.e., the highest variable occurs in the clause is $x_{i}$.
  • $\bullet$
    The trail $M$ is the list of literals evaluated as true (either decided or propagated). The literals in $M$ are organized levels, where each level $i$ consists of a list $M_{i}$ of literals assigned at that level, and the assignment of $x_{i}$. Thus, $M$ has the structure $M=[M_{1};x_{1}\rightarrow assignment[x_{1}];M_{2};x_{2}\rightarrow assignment[x_{2}];...;M_{i};x_{i}\rightarrow assignment[x_{i}];...]$. Each $M_{i}$ contains decided literals (e.g.$\ell_{j}$ if $\ell_{j}$ is decided to true, or $\neg\ell_{j}$ if false), and propagated literals (e.g.$c\vdash\ell_{j}$ if clause $c$ implies $\ell_{j}$ is true, or clause $c$ implies $\neg\ell_{j}$ is false). For example, $M_{i}=[\ell_{1},\ell_{2},c\vdash\ell_{3},\ell_{4},\ldots,\ell_{k}]$.
  • $\bullet$
    $\mathop{\mathtt{value}}(CS_{i},M)$ is the value (true/false/undef) of the conjunction of clauses in $CS_{i}$ by replacing literals in $M$ to be true, and negation of literals in $M$ to be false. Similarly, $\mathop{\mathtt{value}}(c,M)$ is the value of the clause $c$, and $\mathop{\mathtt{value}}(l,M)$ is the value of the literal $l$.
  • $\bullet$
    $\texttt{Solve}(M)$ is the solution interval of $x_{level}$ restricting literals in $M$ to be true, where $M$ has $level$ levels. (Note that assignments of $x_{1},\ldots,x_{level-1}$ are fixed by $M$, so the solution is derived w.r.t. $x_{level}$).
  • $\bullet$
    $\texttt{Consistent}(\ell,M)$ is the consistency (true/false) of the literal $\ell$ and literals in $M$, i.e., whether $x_{level}$ has a solution under the constraints if $\ell$ is true and literals in $M$ are true.
  • $\bullet$
    The “minimal conflicting core of $M$ and $\ell$ on the level $level$” refers to a minimal subset of literals on level $level$ in $M$ which is inconsistent with $\ell$.
  • $\bullet$
    The function $\texttt{explain}(minCore,a:=(assignment[x_{1}],\ldots,assignment[x_{level-1}]))$ constructs a cell (by the single-cell projection operator $\mathop{\mathtt{Proj}}$) of the set of polynomials in $minCore$, denoted by $P\subseteq\mathbb{Q}[x_{1},\ldots,x_{level}]$, and the sample point $a\in\mathbb{R}^{level-1}$. Formally, $\texttt{explain}(minCore,a)=\mathop{\mathtt{Proj}}(P,x_{level},a)$ (see Definition 2.4).
  • $\bullet$
    $\texttt{resolve}(c,M,level)$ is the resolution of the clause $c$ and the conjunction of literals in $M$ on the level $level$.
Next, we explain the our implementation of MCSAT. MCSAT combines model constructing with conflict driven clause learning. The core idea of MCSAT is levelwise constructing a model by assigning variables according to a fixed order, and generating lemmas when encountering conflicts to avoid redundant search. The input of MCSAT is a polynomial formula $F$ with variables $x_{1},\ldots,x_{n}$, such that the relational operator of every atom is ‘$<$’ or ‘$>$’. The output of MCSAT is ‘SAT’ if $F$ is satisfiable, or ‘UNSAT’ if $F$ is unsatisfiable. MCSAT has the following four core strategies.
Assign (lines 2-2): If clauses in $CS_{level}$ are evaluated to be true, then the variable $x_{level}$ is assigned to be one element in the interval $\texttt{Solve}(M)$, and MCSAT’s search enters the next level. If all variables are assigned, MCSAT returns ‘SAT’.
Status Update (lines 2-2, 2-2): The status of MCSAT $mcstatus$ is updated after deciding the value of an undefined literal in a clause (‘Decide’), propagating an undefined literal in a clause (‘Propagate’), or determining $F$ is unsatisfiable if the value of a clause is false (‘UNSAT’).
Consistency Check (black lines in lines 2-2): The decided or propagated literal is checked if it is consistent with $M$. If so, the literal is appended to $M$. Otherwise, MCSAT learns a lemma $lemma$ which is the negation of the conjunction of three components. The first component $cell$ is a set of unsatisfied assignments if literals in $minCore$ and the literal $\ell$ are satisfied; the second component $\neg(\land_{\ell^{\prime}\in minCore}\ell^{\prime})$ consists of literals in $minCore$; the third component is the literal $\ell$. This lemma represents the three components cannot hold simultaneously. Then, $lemma$ is added to $CS_{level}$, and $lemma\rightarrow\neg\ell$ is appended to $M$. If the literal $\ell$ inconsistent with $M$ stems from propagation, MCSAT updates its status $mcstatus$ to ‘UNSAT’.
Conflict Resolve (black lines in lines 2-2): When the status of MCSAT is ‘UNSAT’, MCSAT learns a lemma $lemma$ by resolution. If $lemma$ is empty, MCSAT returns ‘UNSAT’. MCSAT backtracks by canceling the last decided literal. If $x_{level}$ appears in $lemma$, MCSAT backtracks to the last decided literal $\ell^{*}$ in $M$ that is an atom of $lemma$; otherwise, MCSAT backtracks to the last variable assignment in $lemma$.
Overall, the MCSAT algorithm operates through four core strategies working in coordination. Beginning with Assign that assigns the variable when all clauses at the current level evaluate to true, the algorithm may terminate with ‘SAT’ if all variables are successfully assigned. When assignment fails, Status Update yields either ‘Decide’, ‘Propagate’, or ‘UNSAT’ statuses. For ‘Decide’ or ‘Propagate’ statuses, Consistency Check verifies the consistency of the decided or propagated literal with the trail $M$, and any detected inconsistency triggers the lemma learning that involves the single-cell projection operator ($\mathop{\mathtt{Proj}}$, see Definition 2.4). Besides, inconsistency under the status ‘Propagate’ leads to the status ‘UNSAT’ by updating in Status Update. For ‘UNSAT’ statuses, Conflict Resolve conducts lemma learning (an empty lemma makes the algorithm terminate with ‘UNSAT’), and then executes non-chronological backtracking driven by the lemma. This process repeats iteratively until termination.
Input: $F$, a polynomial formula with variables $x_{1},\ldots,x_{n}$ such that the relational operator of every atom is ‘$<$’ or ‘$>$’; $MaxRestart_{1}$, $MaxRestart_{2}$, $MaxJump$, $m$ and $max\_numFailCells$, five positive integers
Output: ‘SAT’ or ‘UNSAT’
1
2 $(status,\alpha,numJump)\leftarrow 2d\text{-LS}(F,MaxRestart_{1},MaxJump,m)$
3 if $status=$‘SAT’ then return ‘SAT’
4
5 $assignment\leftarrow\alpha$, $numFailCells\leftarrow numJump$, $numTLE\leftarrow 0$
[…]
7for $i$ from $2$ to $n$ do $CS_{i}\leftarrow\{c\in\mathop{\mathtt{clauses}}(F)\mid\text{for every }\ell\in\mathop{\mathtt{atoms}}(c),~{}\mathop{\mathtt{poly}}(\ell)\in\mathbb{Q}[x_{1},\ldots,x_{i}]\}\setminus\cup_{j=1}^{i-1}CS_{j}$
8
[…]
10while $\mathop{\mathtt{true}}$ do
11       if $\mathop{\mathtt{value}}(CS_{level},M)=\mathop{\mathtt{true}}$ then
12             if $assignment[x_{level}]$ not in $\textnormal{{Solve}}(M)$ then
13                   $assignment[x_{level}]\leftarrow$ one element in $\textnormal{{Solve}}(M)$
14            $M\leftarrow[M;x_{level}\mapsto assignment[x_{level}]]$
15            $level++$
16            $maxlevel\leftarrow\max(level,maxlevel)$
17            if $level>n$ then return ‘SAT’
18             if $n-2>level>\min(0,4*n,0.9*maxlevel)$ and $numTLE<3$ then
19                   $(status,\alpha,numJump)\leftarrow\allowbreak 2d\text{-LS}(F\mid_{x_{1}=assignment[x_{1}],\ldots,x_{level-1}=assignment[x_{level-1}]},MaxRestart_{2},MaxJump,m)$
20                  if $status=$‘SAT’ then return ‘SAT’
21                   else
22                         for i from $level$ to $n$ do $assignment[x_{i}]\leftarrow\alpha[x_{i}]$
23
24                        $numFailCells\leftarrow numFailCells+numJump$
25
26
27       else
28             if there exists a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{false}}$ then  $mcstatus\leftarrow(\text{`UNSAT'},c)$
29             else if there exists a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{undef}}$ and only one literal $\ell$ in clause $c$ such that $\mathop{\mathtt{value}}(\ell,M)=\mathop{\mathtt{undef}}$ then  $mcstatus\leftarrow(\text{`Propagate'},\ell,c)$
30             else
31                   choose a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{undef}}$
32                  choose a literal $\ell$ in $c$ such that $\mathop{\mathtt{value}}(\ell,M)=\mathop{\mathtt{undef}}$
33                  $mcstatus\leftarrow(\text{`Decide'},\ell,c)$
34
35            if $mcstatus[1]\neq$ ‘UNSAT’  then
36                   $\ell\leftarrow mcstatus$[2], $c\leftarrow mcstatus$[3]
37                  if $\mathop{\mathtt{Consistent}}(\ell,M)$  then
38                         if $status[1]=$ ‘Decide’ then  $M\leftarrow[M;\ell]$
39                         else if $status[1]=$ ‘Propagate’ then  $M\leftarrow[M;c\to\ell]$
40
41                   else
42                         $minCore\leftarrow$ minimal conflicting core of $M$ and $\ell$ on the level $level$
43                        $cell\leftarrow\mathop{\mathtt{explain}}(minCore,(assignment[x_{1}],\ldots,assignment[x_{level-1}]))$
44                        $lemma\leftarrow\neg\left(cell\land(\bigwedge_{\ell^{\prime}\in minCore}\ell^{\prime})\land\ell\right)$
45                        $learnClauses\leftarrow learnClauses\cup\{lemma\}$
46                        $CS_{level}\leftarrow CS_{level}\cup\{lemma\}$
47                        $M\leftarrow[M;lemma\to\neg\ell]$
48                        if $status[1]=$ ‘Propagate’ then
49                               $numFailCells\leftarrow numFailCells+1$
[…]
51
52
53
54            if $mcstatus[1]=$ ‘UNSAT’  then
55                   $c\leftarrow mcstatus[2]$, $lemma\leftarrow\mathop{\mathtt{resolve}}(c,M,level)$
56                  if $lemma$ is empty then return ‘UNSAT’
57
58                  $learnClauses\leftarrow learnClauses\cup\{lemma\}$
59                  if $x_{level}$ appears in $lemma$ then
60                         $CS_{level}\leftarrow CS_{level}\cup\{lemma\}$
61                        $\ell^{*}\leftarrow$ the last decided literal in $M$ satisfying $\ell^{*}\in\mathop{\mathtt{atoms}}(lemma)$
62                        delete the decided literal $\ell^{*}$ and all subsequent terms from $M$
63
64                  else
65                         $tmpLevel\leftarrow$ maximal $i$ of $x_{i}$ appearing in $lemma$
66                        $CS_{tmpLevel}\leftarrow CS_{tmpLevel}\cup\{lemma\}$
67                        delete the assignment $x_{tmpLevel}\mapsto assignment[x_{tmpLevel}]$ and all subsequent terms from $M$
68                        $level\leftarrow tmpLevel$
69
70
71       if $numFailCells>max\_numFailCells$ then
72            return OpenCAD$(F,learnClauses)$
73
Algorithm 2 The Hybrid Framework
5. Experiments
5.1. Experiments Setups
5.1.1. Environment
In this paper, all experiments are conducted on 8-Core 11th Gen Intel(R) Core(TM) i7-11700@2.50GHz with 16GB of memory under the operating system Ubuntu 24.04 (64-bit). Each solver has only one chance to solve each instance within 1200 seconds.
5.1.2. Benchmarks
  • $\bullet$
    SMTLIB: This refers to a filtered subset of the QF_NRA benchmark from the SMT-LIB [^1] http://smtlib.cs.uiowa.edu/ standard benchmark library, containing all instances with only strict inequalities. This selection aligns with our hybrid framework’s input requirements (see Alg. 2). There are 2050 instances in this bechmark.
  • $\bullet$
    Random Instances: Random instances are generated by the random polynomial formula generating function $\mathbf{rf}$ and parameters in (Li et al., 2023, Sect. 7.2), i.e.$\mathbf{rf}$({30,40}, {60,80},{20,30},{10,20},{20,30},{40,60},{3,5}). These instances are almost always satisfiable. Unlike SMTLIB, which has many unit clauses and linear polynomial constraints, the random instances are more complicated in Boolean structure (i.e., they have at least 3 atoms in every clause) and highly nonlinear (i.e., the degree of every polynomial is at least 20). There are 100 instances in this benchmark.
  • $\bullet$
    Specific Instances: Specific instances are all instances in (Li and Xia, 2020, Sect. 5), including Han_$\_n$, P, Hong_$\_n$, Hong2_$\_n$ and C_$\_n\_r$. These instances have particular mathematical properties that are difficult for solvers. There are 21 instances in this benchmark.
5.1.3. Implementation
Algorithm 2 has been implemented as a hybrid solver HELMS [^2] The solvers are avaliable on github. For reasons of anonymity, we cannot provide the address at this time. , and the sample-cell projection operation (Li and Xia, 2020) has been implemented in the MCSAT solver LiMbS [^2] The solvers are avaliable on github. For reasons of anonymity, we cannot provide the address at this time. with Mathematica 14. Besides, we deploy a rollback mechanism on the satisfiability check in $2d$-LS. When an assignment is checked for unsatisfiability, $2d$-LS does a celljump. If the assignment after this cell-jump is satisfied, then $2d$-LS returns ‘SAT’; otherwise, $2d$-LS rolls back to the original assignment.
5.2. Competitors
  • $\bullet$
    Our Main Solver: HELMS is the hybrid solver for $2d$-LS and MCSAT according to the hybrid framework in Algortithm 2.
  • $\bullet$
    Base Solvers: LS is a base local search solver for HELMS and LiMbS is a base MCSAT solver for HELMS.
  • $\bullet$
    SOTA Solvers: Four SOTA solvers from recent SMT Competitions (SMT-COMP) are Z3 (version 4.13.4), CVC5 (version 1.2.0), MathSAT5 (version 5.6.11) and Yices2 (version 2.6.5).
5.2.1. Indicators
#SAT is the number of satisfiable instances and #UNSAT is the number of unsatisfiable instances. #INST is the number of instances. #ALL is #SAT + #UNSAT.
Table 1. The number of instances solved by HELMS, base solvers and SOTA solvers on three benchmarks.
Benchmark #INST HELMS LS LiMbS Z3 CVC5 MathSAT5 Yices2
SMTLIB #SAT 1503 1503 1472 1502 1503 1482 812 1496
#UNSAT 547 540 0 538 536 535 315 538
#ALL 2050 2043 1472 2040 2039 2017 1127 2034
Rand Inst. #SAT 100 100 100 71 16 0 1 23
Spec Inst. #SAT 7 7 4 7 7 4 1 3
#UNSAT 14 14 0 14 5 4 2 3
#ALL 21 21 4 21 12 8 3 6
Total #SAT 1610 1610 1576 1580 1526 1486 814 1522
#UNSAT 561 554 0 552 541 539 317 541
#ALL 2171 2164 0 2132 2067 2025 1131 2063
5.3. Comparison to SOTA Solvers
On the benchmark of SMTLIB, in terms of the number of solved instances, as shown in Table 1, HELMS solves 2043 out of 2050 instances, outperforming SOTA solvers. HELMS is 10, 100 or even 1000 times faster than SOTA solvers on many instances. In Figure 4, we compare the runtime of HELMS and every SOTA solvers for each instance. Our strategy exhibits a bias towards SAT instances, which consequently results in relatively slower performance on UNSAT instances. It is evident that HELMS demonstrates a time advantage on SAT instances.
On the benchmark of random instances, HELMS solves all of 100 instances. The SOTA solver that solves most random instances is Yices2 which only solves 23 instances. CVC5 even solves none of them. Therefore, HELMS outperforms SOTA solvers when dealing with more complicated Boolean structures and nonlinearity.
On the benchmark of specific instances, HELMS successfully solves all 21 instances, whereas SOTA solvers experience timeouts on the majority of both SAT and UNSAT instances. Consequently, HELMS demonstrates superior performance on instances whose mathematical properties pose significant challenges for SOTA solvers.
Overall, HELMS is competitive with SOTA solvers. The cumulative distribution functions (CDFs) of all benchmarks in Figure 6 illustrate that HELMS not only solves the most instances compared to SOTA solvers, but also competitve in runtime. Although the CDF of MathSAT5 lies to the left of HELMS, its number of solved instances is significantly lower compared to that of HELMS. The other three SOTA solvers are positioned to the right of HELMS’s CDF over a relatively extended time range, indicating that HELMS achieves faster solving times across the majority of instances. Especially, the advantage of HELMS on SAT instances is notable. HELMS solves all SAT instances within 2 seconds. The CDFs for SAT instances within 2 seconds are shown in the Figure 5(b), demonstrating the superior performance of HELMS on SAT instances.
(a) Compared to Z3
(b) Compared to CVC5
(c) Compared to MathSAT5
(d) Compared to Yices2
Figure 4. Scatter plots that compare runtime of two pairs of solvers on every instance in all benchmarks. The orange (resp., blue) nodes denote SAT (resp., UNSAT) instances.
5.4. Effectiveness of Proposed Strategies
5.4.1. Effectiveness of $2d$-LS
To show the effectiveness of extending local search with MCSAT, we compare the runtime of $2d$-LS and LS on random instances, which is the most difficult benckmark for SAT instances among the three benckmarks. In Figure 6, all nodes are located above the diagonal and most nodes are near to the left, indicating $2d$-LS improves significantly over LS.
(a) All instances.
(b) SAT instances.
Figure 5. CDFs of HELMS and SOTA solvers on the SMTLIB and all benckmarks. The position of the CDF curve to the left signifies a more rapid solver performance, while a higher curve elevation indicates a greater number of instances that the solver is capable of solving.
Figure 6. Runtime of $2d$-LS and LS on all benchmarks. Points above the diagonal denote instances where $2d$-LS is faster, and points below the diagonal denote instances where LS is faster.
5.4.2. Effectiveness of the Hybrid Framework
Table 2. Stages that HELMS ends in solving instances on all benchmarks.
Benchmark #INST HELMS $2d$-LS MCSAT-2dLS OpenCAD
SMTLIB #SAT 1503 1503 1497 6 0
#UNSAT 547 540 0 535 5
#ALL 2050 2043 1497 541 5
Random Inst. #SAT 100 100 100 0 0
Spec Inst. #SAT 7 7 7 0 0
#UNSAT 14 14 0 11 3
#ALL 21 21 7 11 3
Total #SAT 1610 1610 1604 6 0
#UNSAT 561 554 0 546 8
#ALL 2171 2164 1604 552 8
As shown in Table 1, HELMS effectively integrates the strengths of both LS and LiMbS across instances with diverse characteristics, and outperforms each of them individually. On the benchmark of SMTLIB, the HELMS improves LiMbS on SAT instances owing to $2d$-LS for SAT, and improves on UNSAT instances owing to OpenCAD. LS and LiMbS excel on random and specific instances, respectively, while HELMS combines their strengths. The data presented in Table 2 demonstrates that HELMS primarily relies on $2d$-LS to solve SAT instances and on MCSAT-2dLS to address UNSAT instances. OpenCAD provides additional support for UNSAT instances that MCSAT cannot solve, thereby enabling HELMS to achieve a higher number of solved UNSAT instances compared to MCSAT alone. On both random and specific instances, HELMS effectively combines the strengths of the base solvers, enabling it to successfully solve instances characterized by highly nonlinear constraints, more complicated Boolean structures, and particular mathematical properties. This integration allows HELMS to achieve superior performance across a diverse range of challenging problem types.
6. Conclusion
In this paper, we have advanced local search for SMT-NRA. First, we introduced a two-dimensional cell-jump operation, termed $2d$-cell-jump, which generalizes the key cell-jump operation in existing local search methods for SMT-NRA. Building on this, we proposed an extended local search framework, named $2d$-LS, which integrates MCSAT to realize $2d$-cell-jump in local search. To further improve MCSAT, we implemented the solver LiMbS that utilizes a recently proposed technique called the sample-cell projection operator, which is well suited for CDCL-style search in the real domain and helps guide the search away from conflicting states. Finally, we presented a hybrid framework that exploits the complementary strengths of MCSAT, $2d$-LS and OpenCAD. In this hybrid framework, MCSAT drives $2d$-LS to accelerate the search for a model, $2d$-LS helps MCSAT identify unsatisfiable cells, and OpenCAD is utilized to handle unsatisfiable formulas dominated by algebraic conflicts. We implemented our hybrid framework in the solver HELMS. Experimental results demonstrate that HELMS is competitive with SOTA solvers on the standard benchmark SMTLIB. Moreover, HELMS is superior to SOTA solvers on other benchmarks that have highly nonlinear constraints, more complicated Boolean structure, and particular mathematical properties. The advantages of HELMS over base solvers validate the effectiveness of the proposed methods.