### v2

# A Hybrid SMT-NRA Solver: Integrating 2D Cell-Jump-Based Local Search, MCSAT and OpenCAD

Abstract.
In this paper, we propose a hybrid framework for Satisfiability Modulo the Theory of Nonlinear Real Arithmetic (SMT-NRA for short). First, we introduce a two-dimensional cell-jump move, called $2d$-cell-jump, generalizing the key operation, cell-jump, of the local search method for SMT-NRA. Then, we propose an extended local search framework, named $2d$-LS (following the local search framework, LS, for SMT-NRA), integrating the model constructing satisfiability calculus (MCSAT) framework to improve search efficiency. To further improve the efficiency of MCSAT, we implement a recently proposed technique called sample-cell projection operator for MCSAT, which is well suited for CDCL-style search in the real domain and helps guide the search away from conflicting states. Finally, we present a hybrid framework for SMT-NRA integrating MCSAT, $2d$-LS and OpenCAD, to improve search efficiency through information exchange. The experimental results demonstrate improvements in local search performance, highlighting the effectiveness of the proposed methods.
SMT-NRA, Local Search, MCSAT, OpenCAD, Hybrid Methods, Satisfiability
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Theory of computation Automated reasoningccs: Mathematics of computing Solvers
1. Introduction
Satisfiability Modulo Theories (SMT) is concerned with determining the satisfiability of first-order logic formulas under background theories, such as integer arithmetic, real arithmetic, arrays, bit vectors, strings, and others. This paper concentrates on SMT problems over the theory of quantifier-free nonlinear real arithmetic (NRA), referred to as SMT-NRA. The goal is to determine the satisfiability of polynomial formulas, which are expressed in the form of $\bigwedge_{i}\bigvee_{j}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0,\text{where}~{}\rhd_{ij}\in\{<,>,\leq,\geq,=,\neq\}$ and $p_{ij}(\bar{\boldsymbol{x}})$ are polynomials. SMT-NRA has found widespread applications in various fields, including for example control theory for system verification (Alur, 2011; Cimatti et al., 2013b; Chen and He, 2018), robotics for motion planning and trajectory optimization (Nedunuri et al., 2014; Shoukry et al., 2016; Imeson and Smith, 2019), and software/hardware verification to ensure timing and performance constraints in embedded systems (Trindade and Cordeiro, 2016; Beyer et al., 2018; Li and Gopalakrishnan, 2010; Faure-Gignoux et al., 2024). It also plays a critical role in optimization (Sebastiani and Tomasi, 2012; Li et al., 2014; Bjørner et al., 2015), where nonlinear constraints are common.
Tarski proposed an algorithm in 1951 (Tarski, 1998), solving the problem of quantifier elimination (QE) of the first-order theory over real closed fields, which takes SMT-NRA as a special case. Cylindrical Algebraic Decomposition (CAD), another real QE method introduced by Collins in 1975 (Collins, 1975), can solve polynomial constraints by decomposing space into finitely many regions (called cells) arranged cylindrically. CAD provides a more practical approach to quantifier elimination than Tarski’s procedure though it remains of doubly exponential complexity. In practice, the Model-Constructing Satisfiability Calculus (MCSAT) (de Moura and Jovanović, 2013) is a widely used complete SMT algorithms. MCSAT integrates two solvers from the classical framework into one solver that simultaneously searches for models in both the Boolean structure and the theory structure, thereby constructing consistent Boolean assignments and theory assignments. Several state-of-the-art (SOTA) SMT solvers supporting NRA have been developed over the past two decades. Representative SOTA solvers include Z3 (de Moura and Bjørner, 2008) and Yices2 (Dutertre, 2014), which implement MCSAT, as well as CVC5 (Barbosa et al., 2022) and MathSAT5 (Cimatti et al., 2013a), which use alternative techniques.
The computational complexity of SMT-NRA remains a challenge, motivating research of incomplete solvers that are usually more efficient in finding SAT assignments. Local search, a popular paradigm, has been developed in recent years for SMT-NRA (Li et al., 2023; Li and Cai, 2023; Wang et al., 2024). Local search begins with a theory assignment and approaches a model of the polynomial formula iteratively by moving locally. This process ends when a model is found or other termination conditions are met. The most effective move for real-space search is the ‘cell-jump’ proposed by Li et al. (Li et al., 2023), which leads sample points to different CAD cells via one-dimensional moves.
The complementarity between complete methods and local search has led to the development of hybrid solvers for related problems. Notable examples include the hybridization of Conflict-Driven Clause Learning (CDCL) and local search for SAT (Letombe and Marques-Silva, 2008; Zhang et al., 2024), as well as the combination of CDCL(T) and local search (Zhang et al., 2024) for solving satisfiability modulo the theory of nonlinear integer arithmetic, SMT-NIA for short. These studies suggest that a similar hybrid approach may hold promise for solving SMT-NRA.
In this paper, we aim at advancing local search in SMT-NRA on problems in the form of
$\bigwedge_{i}\bigvee_{j}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0,\text{where}~{}\rhd_{ij}\in\{<,>,\neq\}~{}\text{and}~{}p_{ij}\in\mathbb{Q}[\bar{\boldsymbol{x}}],$
through the integration of MCSAT and make the following contributions:
  • $\bullet$
    We propose a new cell-jump mechanism, called $2d$-cell-jump, which supports two-dimensional search and may find models faster.
  • $\bullet$
    We propose an extended local search framework, named $2d$-LS, integrating the MCSAT framework to improve search efficiency.
  • $\bullet$
    Inspired by the work (Zhang et al., 2024) of Zhang et al. on the combination of CDCL(T) and local search for SMT-NIA, we design a hybrid framework for SMT-NRA that exploits the complementary strengths of MCSAT, $2d$-LS and OpenCAD (Han et al., 2014). In this framework, $2d$-LS accelerates model search within MCSAT and assists in identifying unsatisfiable cells, which are then used as a heuristic signal to trigger a switch to OpenCAD for handling unsatisfiable formulas dominated by algebraic conflicts.
  • $\bullet$
    The above proposed methods have been implemented as a solver called HELMS. When implementing MCSAT, we use a recently proposed technique called sample-cell projection operator for MCSAT, which further improves the efficiency of MCSAT. Comparison to SOTA solvers on a large number of benchmarks shows that the newly proposed methods are effective.
The rest of this paper is organized as follows. Sect. 2 introduces preliminaries of the problem in SMT-NRA, the local search solver and the sample-cell projection for MCSAT. Sect. 3 extends local search to $2d$-LS, introducing the new cell-jump. Sect. 4 outlines the hybrid method that combines $2d$-LS, MCSAT and OpenCAD. The experimental results in Sect. 5 demonstrate that our hybrid SMT-NRA solver is efficient and integrates complementary strengths of $2d$-LS, MCSAT and OpenCAD. Finally, Sect. 6 concludes this paper.
2. Preliminaries
2.1. Problem Statement
Let $\bar{\boldsymbol{x}}=(x_{1},\ldots,x_{n})$ be a vector of variables. Denote by $\mathbb{Q}$, $\mathbb{R}$, $\mathbb{Z}$ and $\mathbb{N}$ the set of rational numbers, real numbers, integers and natural numbers, respectively. Let $\mathbb{Q}[\bar{\boldsymbol{x}}]$ be the ring of polynomials in the variables $x_{1},\ldots,x_{n}$ with coefficients in $\mathbb{Q}$.
Definition 2.1 (Polynomial Formula).
The following formula
(1) $F=\bigwedge_{i=1}^{M}\bigvee_{j=1}^{m_{i}}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$
is called a polynomial formula, where $1\leq i\leq M<+\infty,1\leq j\leq m_{i}<+\infty$, $p_{ij}\in\mathbb{Q}[\bar{\boldsymbol{x}}]$ and $\rhd_{ij}\in\{<,>,\leq,\geq,=,\neq\}$. Moreover, $\bigvee_{j=1}^{m_{i}}p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$ is called a polynomial clause or simply a clause, and $p_{ij}(\bar{\boldsymbol{x}})\rhd_{ij}0$ is called an atomic polynomial formula or simply an atom. As is customary, we call an atomic polynomial formula or its negation a polynomial literal or simply a literal.
For any polynomial formula $F$, a complete assignment is a mapping $\alpha:\bar{\boldsymbol{x}}\to\mathbb{R}^{n}$ such that $x_{1}\mapsto a_{1},\ldots,x_{n}\mapsto a_{n}$, where every $a_{i}\in\mathbb{R}$. We denote by $\alpha[x_{i}]$ the assigned value $a_{i}$ of the variable $x_{i}$. With slight abuse of notation, we sometimes represent a complete assignment simply by the real vector $(a_{1},\ldots,a_{n})$. An atom is true under $\alpha$ if it evaluates to true under $\alpha$, and otherwise it is false under $\alpha$, A clause is satisfied under $\alpha$ if at least one atom in the clause is true under $\alpha$, and falsified under $\alpha$ otherwise. When the context is clear, we simply say a true (or false) atom and a satisfied (or falsified) clause. A polynomial formula is satisfiable (SAT) if there exists a complete assignment in $\mathbb{R}^{n}$ such that all clauses in the formula are satisfied, and such an assignment is a model to the polynomial formula. A polynomial formula is unsatisfiable (UNSAT) if any assignment is not a model.
Example 2.2.
(A running example) We take the following polynomial formula as a running example
$F_{r}=f_{1,r}<0\land f_{2,r}<0,$
where $r\in\mathbb{N}$, $f_{1,r}=x^{2}+y_{1}^{2}+\cdots+y_{r}^{2}-z^{2}$ and $f_{2,r}=(x-3)^{2}+y_{1}^{2}+\cdots+y_{r}^{2}+z^{2}-5$. Let $\ell_{1,r}$ denote atom $f_{1,r}<0$ and $\ell_{2,r}$ denote atom $f_{2,r}<0$. Under the assignment $(x,y_{1},\ldots,y_{r},z)\mapsto(\frac{3}{2},0,\ldots,0,\frac{8}{5})$, both $\ell_{1,r}$ and $\ell_{2,r}$ are true, and thus $F_{r}$ is satisfiable.
The problem we consider in this paper is to determine the satisfiability of polynomial formulas in the form of (1) with $\rhd_{ij}\in\{<,>,\neq\}.$
2.2. Cell-Jump Operation in Local Search
Li et al. propose a local search algorithm (Li et al., 2023, Alg. 3) for solving SMT-NRA. The key point of the algorithm is the cell-jump operation (Li et al., 2023, Def. 11 & Alg. 2), which updates the current assignment along a straight line with given direction.
  • $\bullet$
    Cell-Jump Along a Coordinate Axis Direction: The first type of cell-jump moves along one coordinate axis direction to update the current assignment. For instance, consider an SMT-NRA problem with two variables $x$ and $y$. Note that the search space is $\mathbb{R}^{2}$. This type of cell-jump moves either along the $x$-axis direction, i.e., updating the assigned value of the first variable $x$, or along the $y$-axis direction, i.e., updating the assigned value of the second variable $y$.
  • $\bullet$
    Cell-Jump Along a Given Direction: The second type of cell-jump moves along any given direction. For instance, consider an SMT-NRA problem with two variables. Given a straight line with the direction $(3,4)$, one cell-jump moves from assignment $(a_{1},a_{2})$ to a new assignment $(a_{1}+3t,a_{2}+4t)$ along the line. Such movement enables a more comprehensive exploration of the search space, potentially facilitating the rapid discovery of solutions.
2.3. Sample-Cell Projection Operator for MCSAT
In our MCSAT implementation, we use a projection operator, called sample-cell projection operator proposed in (Li and Xia, 2020), which is essentially the same as the ‘biggest cell’ heuristic in the recent work (Nalbach et al., 2024). Compared with (Brown, 2013), the sample-cell projection operator is better suited for integration with the MCSAT framework, enabling both efficient solving and lazy evaluation. The sample-cell projection operator aims at generating a larger cell (which is a region). The impact on MCSAT is that, for unsatisfiable problems, excluding a larger unsatisfiable cell generally leads to higher efficiency in subsequent CDCL-style search. We introduce the sample-cell projection operator in this subsection.
Let $f,g\in\mathbb{Q}[\bar{\boldsymbol{x}}]$, $F$ be a finite subset of $\mathbb{Q}[\bar{\boldsymbol{x}}]$ and $a\in\mathbb{R}^{n}$. Denote by $\mathop{\mathtt{disc}}(f,x_{i})$ and $\mathop{\mathtt{res}}(f,g,x_{i})$ the discriminant of $f$ with respect to $x_{i}$ and the resultant of $f$ and $g$ with respect to $x_{i}$, respectively. The order of $f$ at $a$ is defined as
${\mathop{\mathtt{order}}}_{a}(f)=\min($$\{k\in\mathbb{N}\mid\text{some partial derivative of total order }k\text{ of }f$
[…]
We call $f$ order-invariant on $S\subseteq\mathbb{R}^{n}$, if ${\mathop{\mathtt{order}}}_{a_{1}}(f)={\mathop{\mathtt{order}}}_{a_{2}}(f)$ for any $a_{1},a_{2}\in S$. Note that order-invariance implies sign-invariance. We call $F$ a square-free basis in $\mathbb{Q}[\bar{\boldsymbol{x}}]$, if the elements in $F$ are of positive degrees, primitive, square-free and pairwise relatively prime.
Definition 2.3 (Analytic Delineability).
(Collins, 1975) Let $r\geq 1$, $S$ be a connected sub-manifold of $\mathbb{R}^{r-1}$ and $f\in\mathbb{Q}[x_{1},\ldots,x_{r}]\setminus\{0\}$. The polynomial $f$ is called analytic delineable on $S$, if there exist finitely many analytic functions $\theta_{1},\ldots,\theta_{k}:S\rightarrow\mathbb{R}$ (for $k\geq 0$) such that
  • $\bullet$
    $\theta_{1}<\cdots<\theta_{k}$,
  • $\bullet$
    the set of real roots of the univariate polynomial $f(a,x_{r})$ is $\{\theta_{1}(a),\ldots,\theta_{k}(a)\}$ for all $a\in S$, and
  • $\bullet$
    there exist positive integers $m_{1},\ldots,m_{k}$ such that for any $a\in S$ and for $j=1,\ldots,k$, the multiplicity of the root $\theta_{j}(a)$ of $f(a,x_{r})$ is $m_{j}$.
Let sample point $a=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and $F=\{f_{1},\ldots,f_{s}\}\subseteq\mathbb{Q}[\bar{\boldsymbol{x}}]\setminus\{0\}$. Consider the real roots of univariate polynomials in
(2) $\{f_{1}(a_{1},\ldots,a_{n-1},x_{n}),\ldots,f_{s}(a_{1},\ldots,a_{n-1},x_{n})\}\setminus\{0\}.$
Denote by $\gamma_{i,k}~{}(\in\mathbb{R})$ the $k$-th real root of $f_{i}(a_{1},\ldots,a_{n-1},x_{n})$. We define the sample polynomial set of $F$ at $a$, denoted by $\mathop{\mathtt{s\_poly}}(F,x_{n},a)$, as follows.
  1. (1)
    If there exists $\gamma_{i,k}$ such that $\gamma_{i,k}=a_{n}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i}\}$.
  2. (2)
    If there exist two real roots $\gamma_{i_{1},k_{1}}$ and $\gamma_{i_{2},k_{2}}$ such that $\gamma_{i_{1},k_{1}}<a_{n}<\gamma_{i_{2},k_{2}}$ and the open interval $(\gamma_{i_{1},k_{1}},\gamma_{i_{2},k_{2}})$ contains no $\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i_{1}},f_{i_{2}}\}$.
  3. (3)
    If there exists $\gamma_{i^{\prime},k^{\prime}}$ such that $a_{n}>\gamma_{i^{\prime},k^{\prime}}$ and for all $\gamma_{i,k}$, $\gamma_{i^{\prime},k^{\prime}}\geq\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i^{\prime}}\}$.
  4. (4)
    If there exists $\gamma_{i^{\prime},k^{\prime}}$ such that $a_{n}<\gamma_{i^{\prime},k^{\prime}}$ and for all $\gamma_{i,k}$, $\gamma_{i^{\prime},k^{\prime}}\leq\gamma_{i,k}$, then $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\{f_{i^{\prime}}\}$.
  5. (5)
    Specially, if every polynomial in (2) has no real roots, define $\mathop{\mathtt{s\_poly}}(F,x_{n},a)=\emptyset$.
For every $f\in F$, suppose $f=c_{m}x_{n}^{d_{m}}+c_{m-1}x_{n}^{d{m-1}}+\cdots+c_{0}x_{n}^{d_{0}}$, where every $c_{i}\in\mathbb{Q}[x_{1},\ldots,x_{n-1}]\setminus\{0\}$, $d_{i}\in\mathbb{N}$ and $d_{m}>d_{m-1}>\cdots>d_{0}$. If there exists $j\in\mathbb{N}$ such that $c_{j}(a_{1},\ldots,a_{n-1})\neq 0$ and $c_{j^{\prime}}(a_{1},\ldots,a_{n-1})=0$ for any $j^{\prime}>j$, then define the sample coefficients of $f$ at $a$ as $\mathop{\mathtt{s\_coeff}}(f,x_{n},a)=\{c_{m},c_{m-1},\ldots,c_{j}\}$. Otherwise, define $\mathop{\mathtt{s\_coeff}}(f,x_{n},a)=\{c_{m},c_{m-1},\ldots,c_{j}\}$
Definition 2.4 (Sample-Cell Projection).
(Li and Xia, 2020) Suppose $F$ is a finite polynomial subset of $\mathbb{Q}[\bar{\boldsymbol{x}}]\setminus\{0\}$ and $a=(a_{1},\ldots,a_{n-1})\in\mathbb{R}^{n-1}$. The sample-cell projection of $F$ on $x_{n}$ at $a$ is defined as
$\mathop{\mathtt{Proj}}(F,x_{n},a)=$ $\bigcup_{f\in F}\{\mathop{\mathtt{s\_coeff}}(f,x_{n},a)\}\cup\bigcup_{f\in F}\{\mathop{\mathtt{disc}}(f,x_{n})\}\cup\bigcup_{\begin{subarray}{c}f\in F,\\ g\in\mathop{\mathtt{s\_poly}}(F,x_{n},a),\\ f\neq g\end{subarray}}\{\mathop{\mathtt{res}}(f,g,x_{n})\}.$
We refer to $\mathop{\mathtt{Proj}}$ as the sample-cell projection operator.
The following theorem establishes the correctness of the sample-cell projection, i.e., the cell generated by $\mathop{\mathtt{Proj}}$ is a region containing the given sample point $a$, and each polynomial in the set $F$ is sign-invariant in this region.
Theorem 2.5.
(Li and Xia, 2020) Let $n\geq 2$, $F$ be a square-free basis in $\mathbb{Q}[\bar{\boldsymbol{x}}]$, $a=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$, and $S$ be a connected sub-manifold of $\mathbb{R}^{n-1}$ such that $(a_{1},\ldots,a_{n-1})\in S$. If every element of $\mathop{\mathtt{Proj}}(F,x_{n},a)$ is order-invariant on $S$, then every element of $F$ either vanishes identically on $S$ or is analytic delineable on $S$.
The sample-cell projection operator can be used for explanation in MCSAT, facilitating conflict-driven clause learning. When a conflict occurs, MCSAT can invoke $\mathop{\mathtt{Proj}}$ on the polynomials involved and the current assignment to derive a clause representing an unsatisfiable cell – i.e., a region responsible for the conflict. This clause prunes the search space by excluding the conflicting region from further exploration. Details of this implementation are presented in Sect. 4.2.
3. Extending Local Search with MCSAT
In Sect. 3.1, we propose a new cell-jump operation, called $2d$-cell-jump. Comparing to cell-jump in (Li et al., 2023, Alg. 2), $2d$-cell-jump allows searching for a model in a plane instead of along a line. In Sect. 3.2, based on $2d$-cell-jump, we develop a new local search algorithm for SMT-NRA, called $2d$-LS. The algorithm can be considered as an extension of LS (Li et al., 2023, Alg. 3).
3.1. New Cell-Jump: $2d$-Cell-Jump
Remark that the cell-jump operation proposed in (Li et al., 2023) is limited to searching for a solution along a straight line, which is one-dimensional. With the assistance of MCSAT, the search process can be extended into higher dimensional space. In this subsection, we define $2$-dimensional cell-jump, $2d$-cell-jump for short, expanding the cell-jump move from a line parallel to a coordinate axis to a plane parallel to an axes plane, and from a given line to a given plane. Theoretically, this approach can be generalized to $D$-dimensional space, where $D\geq 2$. To improve the efficiency of MCSAT, we set $D=2$.
3.1.1. New Sample Point
Note that sample points (Li et al., 2023, Def. 10) are candidate assignments to move to in the original cell-jump operation. To define the new cell-jump operation, we first introduce new sample points. These sample points are generated using MCSAT, where we seek one model (if exists) for atomic polynomial formulas by fixing all variables except for two specific ones.
Definition 3.1 (Sample Point).
Let $n\geq 2$. Consider atom $\ell:p(\bar{\boldsymbol{x}})>0$ (or atom $\ell:p(\bar{\boldsymbol{x}})<0$), and suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. For any pair of distinct variables $x_{i}$ and $x_{j}$ $(i<j)$, let $p^{*}(x_{i},x_{j})=p(a_{1},\ldots,a_{i-1},x_{i},a_{i+1},\ldots,a_{j-1},x_{j},a_{j+1},\ldots,a_{n})\allowbreak\in\mathbb{Q}[x_{i},x_{j}]$. If $p^{*}(x_{i},x_{j})>0$ (or $p^{*}(x_{i},x_{j})<0$) is satisfiable and $(b_{i},b_{j})\in\mathbb{Q}^{2}$ is a model of it, then $(a_{1},\ldots,a_{i-1},b_{i},a_{i+1},\ldots,a_{j-1},b_{j},a_{j+1},\ldots,a_{n})$ is a sample point of $\ell$ with respect to (w.r.t.) $x_{i},x_{j}$ under $\alpha$.
Remark that a sample point of an atom is a model of it. In practice, we use MCSAT to determine the satisfiability of $p^{*}(x_{i},x_{j})>0$ (or $p^{*}(x_{i},x_{j})<0$). The reason is that MCSAT can quickly determine the satisfiability of two-variable polynomial formulas. If a formula only contains one variable, the formula can be solved directly by real root isolation.
Example 3.2.
Consider atoms $\ell_{1,r}:f_{1,r}<0$ and $\ell_{2,r}:f_{2,r}<0$ in Example 2.2. Suppose the current assignment is $\alpha:(x,y_{1},\ldots,y_{r},z)\mapsto(0,0,\ldots,0,0)$. Keeping the variables $x$ and $z$, we have $f_{1,r}|\allowbreak_{y_{1}=0,\ldots,y_{r}=0}=x^{2}-z^{2}$, and $(1,2)$ is a model of it. So, $(1,0,\ldots,0,2)$ is a sample point of $\ell_{1,r}$ w.r.t. $x,z$ under $\alpha$. Keeping the variables $y_{r}$ and $z$, we obtain $f_{2,r}|\allowbreak_{x=0,y_{1}=0,\ldots,y_{r-1}=0}=y_{r}^{2}+z^{2}+4$, which has no models. So, there is no sample point of $\ell_{2,r}$ w.r.t. $y_{r},z$ under $\alpha$.
Let $r=1$. We have $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}$, $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5$ and $\alpha:(x,y_{1},z)\mapsto(0,0,0)$. As shown in Fig. 1, the region above the orange cone satisfies $\ell_{1,1}$, and that inside the blue sphere satisfies $\ell_{2,1}$. So, every point in the intersection of the region above the orange cone and the green plane is a sample point of $\ell_{1,1}$ w.r.t. $x,z$ under $\alpha$, such as point $(1,0,2)$. The region inside the blue sphere and the red plane has no intersection. So, there is no sample point of $\ell_{2,1}$ w.r.t. $y_{r},z$ under $\alpha$.
Figure 1. The figure of a sample point/$2d$-cell-jump operation in a plane parallel to an axes plane. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes $\{(x,0,z)\mid x\in\mathbb{R},z\in\mathbb{R}\}$, and the red plane denotes $\{(0,y_{1},z)\mid y_{1}\in\mathbb{R},z\in\mathbb{R}\}$. The current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$ (i.e., the vertex of the yellow cone). Point $(1,0,2)$ is a sample point of $\ell_{1,1}$ w.r.t. $x,z$ under $\alpha$. There is a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{1},e_{3})$ operation jumping from $(0,0,0)$ to $(1,0,2)$ in the green plane.
3.1.2. New Cell-Jump
For any $i$ $(1\leq i\leq n)$, let $e_{i}=(0,\ldots,1,\allowbreak\ldots,0)$ be a vector in $\mathbb{R}^{n}$ with $1$ in the $i$-th position. For any point $\alpha=(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and any two linearly independent vectors $d_{1},d_{2}\in\mathbb{R}^{n}$, let
$\alpha+\langle d_{1},d_{2}\rangle=\{a_{1}e_{1}+\cdots+a_{n}e_{n}+\lambda_{1}d_{1}+\lambda_{2}d_{2}\mid\lambda_{1}\in\mathbb{R},\lambda_{2}\in\mathbb{R}\},$
which is a plane spanned by vectors $d_{1}$ and $d_{2}$, passing through point $\alpha$. Specially, for any $i,j$ $(1\leq i<j\leq n)$, $(0,\ldots,0)+\langle e_{i},e_{j}\rangle$ is called an axes plane. And $\alpha+\langle e_{i},e_{j}\rangle$ is a plane parallel to an axes plane.
Definition 3.3 ($2d$-Cell-Jump in a Plane Parallel to an Axes Plane).
Suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. Let $\ell$ be a false atom $p(\bar{\boldsymbol{x}})<0$ or $p(\bar{\boldsymbol{x}})>0$. For each pair of distinct variables $x_{i}$ and $x_{j}$ $(i<j)$ such that there exists a sample point $\alpha_{s}$ of $\ell$ w.r.t. $x_{i},x_{j}$ under $\alpha$, there exists a $2d$-cell-jump operation in the plane $\alpha+\langle e_{i},e_{j}\rangle$, denoted as ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell,\alpha,e_{i},e_{j})$, updating $\alpha$ to $\alpha_{s}$ (making $\ell$ become true).
Example 3.4.
Consider polynomial formula $F_{r}=f_{1,r}<0\land f_{2,r}<0$ in Example 2.2. Suppose the current assignment is $\alpha:(x,y_{1},\ldots,y_{r},z)\mapsto(0,0,\ldots,0,0)$. Both atoms $\ell_{1,r}:f_{1,r}<0$ and $\ell_{2,r}:f_{2,r}<0$ are false under $\alpha$. Recalling Example 3.2, $(1,0,\ldots,0,2)$ is a sample point of $\ell_{1,r}$ w.r.t. $x,z$ under $\alpha$, and there is no sample point of $\ell_{2,r}$ w.r.t. $y_{r},z$ under $\alpha$. So, there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,r},\alpha,e_{1},e_{r+2})$ operation, updating $\alpha$ to $(1,0,\ldots,0,2)$, and no ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,r},\alpha,e_{r+1},e_{r+2})$ operation exists. Let $r=1$. As shown in Fig. 1, the ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{1},e_{3})$ operation jumps from $(0,0,0)$ to $(1,0,2)$ in the green plane, and there is no ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,e_{2},e_{3})$ operation in the red plane.
Example 3.5.
(a) One cell-jump move.
(b) One $2d$-cell-jump move.
(c) Two cell-jump moves.
(d) Two $2d$-cell-jump moves.
Figure 2. Comparison between cell-jump (Li et al., 2023, Def. 11) and $2d$-cell-jump. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes $\{(x,0,z)\mid x\in\mathbb{R},z\in\mathbb{R}\}$, and the red plane denotes $\{(0,y_{1},z)\mid y_{1}\in\mathbb{R},z\in\mathbb{R}\}$. The vertex of the yellow cone $(0,0,0)$ is the current assignment.
Let $r=1$ and consider polynomial formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$ in Example 2.2, where $\ell_{1,1}$ denotes atom $f_{1,1}<0$ and $\ell_{2,1}$ denotes atom $f_{2,1}<0$. Suppose the current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$. In Fig. 2, we compare the cell-jump operation $\mathop{\mathtt{cjump}}$ along a line parallel to a coordinate axis (Li et al., 2023, Def. 11) and the $2d$-cell-jump operation ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}$ in a plane parallel to an axes plane.
The vertex of the yellow cone $(0,0,0)$ is the current assignment. From point $(0,0,0)$, there exists a $\mathop{\mathtt{cjump}}(z,\ell_{1,1})$ operation jumping to $(0,0,2)$ along the $z$-axis (as shown in Fig. 2), and there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{1,1},\alpha,e_{2},e_{3})$ jumping to the same point in the $y_{1}z$-plane (as shown in Fig. 2). After a one-step move, both cell-jump and $2d$-cell-jump make $\ell_{1,1}$ become true.
Note that $(0,0,2)$ is not a model of $\ell_{2,1}$. Consider cell-jump and $2d$-cell-jump of $\ell_{2,1}$ from point $(0,0,2)$. There exists a $\mathop{\mathtt{cjump}}(z,\ell_{2,1})$ operation jumping to $(\frac{5}{2},0,2)$ along the $x$-axis (as shown in Fig. 2), while there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,e_{1},e_{2})$ jumping to $(\frac{3}{2},0,\frac{8}{5})$ in the $xy$-plane (as shown in Fig. 2). Both the second jumps make $\ell_{2,1}$ become true. It is easy to check that $(\frac{5}{2},0,2)$ is not a model of $\ell_{1,1}$, but $(\frac{3}{2},0,\frac{8}{5})$ is. So, after a two-step move, $2d$-cell-jump finds a model of formula $F_{1}$, while cell-jump does not. The reason is that the $2d$-cell-jump operation searches in a plane, covering a wider search area, potentially leading to faster model discovery.
Definition 3.6 ($2d$-Cell-Jump in a Given Plane).
Suppose the current assignment is $\alpha:(x_{1},\ldots,x_{n})\mapsto(a_{1},\ldots,a_{n})$ where $a_{i}\in\mathbb{Q}$. Let $\ell$ be a false atom $p(\bar{\boldsymbol{x}})<0$ (or $p(\bar{\boldsymbol{x}})>0$). Given two linearly independent vectors $d_{1}=(d_{1,1},\ldots,d_{1,n})$ and $d_{2}=(d_{2,1},\ldots,d_{2,n})$ in $\mathbb{Q}^{n}$, introduce two new variables $t_{1},t_{2}$ and replace every $x_{i}$ with $a_{i}+d_{1,i}t_{1}+d_{2,i}t_{2}$ in $p(\bar{\boldsymbol{x}})$ to obtain a bivariate polynomial $p^{*}(t_{1},t_{2})$. If there exists a sample point $(t_{1}^{*},t_{2}^{*})$ of $p^{*}(t_{1},t_{2})<0$ (or $p^{*}(t_{1},t_{2})>0$) w.r.t. $t_{1},t_{2}$ under assignment $(t_{1},t_{2})\mapsto(0,0)$, then there exists a $2d$-cell-jump operation in the plane $\alpha+\langle d_{1},d_{2}\rangle$, denoted as ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell,\alpha,d_{1},d_{2})$, updating $\alpha$ to $\alpha+t_{1}^{*}d_{1}+t_{2}^{*}d_{2}$ (making $\ell$ become true).
Example 3.7.
Figure 3. The figure of a $2d$-cell-jump operation in a given plane. The graphs of $f_{1,1}=x^{2}+y_{1}^{2}-z^{2}=0$ and $f_{2,1}=(x-3)^{2}+y_{1}^{2}+z^{2}-5=0$ are showed as the orange cone and the blue sphere, respectively. The region above the orange cone satisfies $\ell_{1,1}:f_{1,1}<0$, and that inside the blue sphere satisfies $\ell_{2,1}:f_{2,1}<0$. The green plane denotes the plane $\langle(0,1,0),(15,0,16)\rangle$. The current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$ (i.e., the vertex of the yellow cone). From $(0,0,0)$, there is a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,(0,1,0),(15,0,16))$ operation jumping to $(\frac{3}{2},0,\frac{8}{5})$ in the green plane, which is a model to formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$.
Let $r=1$ and consider polynomial formula $F_{1}=\ell_{1,1}\land\ell_{2,1}$ in Example 2.2, where $\ell_{1,1}$ denotes atom $f_{1,1}<0$ and $\ell_{2,1}$ denotes atom $f_{2,1}<0$. Suppose the current assignment is $\alpha:(x,y_{1},z)\mapsto(0,0,0)$, and $\ell_{2,1}$ is false under $\alpha$. Given two linearly independent vectors $d_{1}=(0,1,0)$ and $d_{2}=(15,0,16)$, consider $2d$-cell-jump operations of $\ell_{2,1}$ in the plane $\alpha+\langle d_{1},d_{2}\rangle$ (the green plane in Fig. 3). Replacing $(x,y_{1},z)$ with $(15t_{2},t_{1},16t_{2})$ in $f_{2,1}$, we get a bivariate polynomial $f_{2,1}^{*}=t_{1}^{2}+481t_{2}^{2}-90t_{2}+4$. It is easy to check that $(0,\frac{1}{10})$ is a sample point of $f_{2,1}^{*}<0$. So, there exists a ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,d_{1},d_{2})$ operation, updating $\alpha$ to $\alpha+t_{1}^{*}d_{1}+t_{2}^{*}d_{2}=(\frac{3}{2},0,\frac{8}{5})$. As shown in Fig. 3, from current location $(0,0,0)$, the ${\mathop{\mathtt{2d}}}{\text{-}}{\mathop{\mathtt{cjump}}}(\ell_{2,1},\alpha,d_{1},d_{2})$ operation jumps to point $(\frac{3}{2},0,\frac{8}{5})$ in the green plane. In fact, $(\frac{3}{2},0,\frac{8}{5})$ is not only a model to $\ell_{2,1}$ but also a model to formula $F_{1}$.
3.2. New Local Search Algorithm: $2d$-LS
Input: $F$, a polynomial formula with variables $x_{1},\ldots,x_{n}$ such that the relational operator of every atom is ‘$<$’ or ‘$>$’; $MaxRestart$, $MaxJump$ and $m$, three positive integers
Output: $status$, ‘SAT’ or ‘UNKNOWN’; $\alpha$, an assignment; $numJump$, a positive integer
1 $numRestart,numJump\leftarrow 1$
2while $numRestart\leq MaxRestart$ do
3       $\alpha\leftarrow(a_{1},\ldots,a_{n})$, an initial complete assignment in $\mathbb{Q}^{n}$
4      while $numJump\leq MaxJump$ do
5             if $F(\alpha)=$True then
6                  return ‘SAT’, $\alpha$, $numJump$
7            if there exists a cell-jump operation (Li et al., 2023, Def. 11) along a line parallel to a coordinate axis with positive score [^1] The scoring function and weighting scheme are from (Li et al., 2023). Their role is to assess candidate sample points. A positive score means that the operation of updating an assignment moves closer to assignments that satisfy the given polynomial formula, with higher scores indicating closer.  then
8                   perform such an operation with the highest score to update $\alpha$
9
10            else
11                   generate $2m$ random vectors $d_{1},\ldots,d_{2m}$, where $d_{i}\in\mathbb{Q}^{n}$
12                  $L\leftarrow\{\alpha+\langle d_{1}\rangle,\ldots,\alpha+\langle d_{2m}\rangle\}$, herein $\alpha+\langle d_{i}\rangle=\{a_{1}e_{1}+\cdots+a_{n}e_{n}+\lambda d_{i}\mid\lambda\in\mathbb{R}\}$ is a random line passing through $\alpha$ with direction $d_{i}$
13                  if there exists a cell-jump operation (Li et al., 2023, Alg. 2) along a line in $L$ with positive score then
14                         perform such an operation with the highest score to update $\alpha$
15
16                  else if there exists a $2d$-cell-jump operation in a plane parallel to an axes plane with positive score then
17                        perform such an operation with the highest score to update $\alpha$
18                   else
19                         $P\leftarrow\{\alpha+\langle d_{1},d_{2}\rangle,\ldots,\alpha+\langle d_{2m-1},d_{2m}\rangle\}$, where every $\alpha+\langle d_{i},d_{j}\rangle$ is a random plane
20                        if there exists a $2d$-cell-jump operation in a plane in $P$ with positive score then
21                              perform such an operation with the highest score to update $\alpha$
22
23                        else
24                              break
25
26
27
28            $numJump$++
29      $numRestart$++
return ‘UNKNOWN’, $\alpha$, $numJump$
Algorithm 1 $2d$-LS
Based on the new cell-jump operation, we develop a new local search algorithm, named $2d$-LS. In fact, $2d$-LS is a generalization of LS (Li et al., 2023, Alg. 3).
Recall that LS adopts a two-step search framework. Building upon LS, $2d$-LS utilizes a four-step search framework, where the first two steps are the same as LS. These four steps are described as follows, also see Alg. 1 for reference.
  1. (1)
    Try to find a cell-jump operation (Li et al., 2023, Def. 11) along a line that passes through the current assignment point with a coordinate axis direction.
  2. (2)
    If the first step fails, generate $2m$ random vectors $d_{1},\ldots,d_{2m}$, where $m\geq 1$. Attempt to perform a cell-jump operation (Li et al., 2023, Alg. 2) along a random line, which passes through the current assignment point with direction $d_{i}$.
  3. (3)
    If the second step fails, try to find a $2d$-cell-jump operation in a plane that passes through the current assignment point and is parallel to an axes plane.
  4. (4)
    If the third step fails, attempt to perform a $2d$-cell-jump operation in a random plane, which is spanned by vectors $d_{i},d_{j}$ and passes through the current assignment point.
Note that cell-jump/$2d$-cell-jump operations are performed on false atoms. False atoms can be found in both falsified clauses and satisfied clauses, where the former are of greater significance in satisfying a polynomial formula. Thus, in every step above, we adopt the two-level heuristic in (Cai et al., 2022, Sect. 4.2). First, try to perform a cell-jump/$2d$-cell-jump operation to make false atoms in falsified clauses become true. If such operation does not exist, then seek one to correct false atoms in satisfiable clauses. (Due to space constraints, this two-level heuristic is not explicitly written in Alg. 1.)
Moreover, there are two noteworthy points in $2d$-LS. First, the algorithm employs the restart mechanism to avoid searching nearby one initial assignment point. The maximum number of restarts $MaxRestart$ is set in the outer loop. Second, in every restart, we limit the number of cell-jumps to avoid situations where there could be infinitely many cell-jumps between two sample points. The maximum number of cell-jumps $MaxJump$ is set in the inner loop.
4. A Hybrid Framework for SMT-NRA
Input: $F$, a polynomial formula with variables $x_{1},\ldots,x_{n}$ such that the relational operator of every atom is ‘$<$’ or ‘$>$’; $MaxRestart_{1}$, $MaxRestart_{2}$, $MaxJump$, $m$ and $max\_numFailCells$, five positive integers
Output: ‘SAT’ or ‘UNSAT’
1
2 $(status,\alpha,numJump)\leftarrow 2d\text{-LS}(F,MaxRestart_{1},MaxJump,m)$
3 if $status=$‘SAT’ then return ‘SAT’
4
5 $assignment\leftarrow\alpha$, $numFailCells\leftarrow numJump$, $maxlevvel\leftarrow 0$
6for $i$ from $1$ to $n$ do  $CS_{i}\leftarrow\{c\in\mathop{\mathtt{clauses}}(F)\mid\mathop{\mathtt{level}}(c)=i\}$
7
[…]
9while $\mathop{\mathtt{true}}$ do
10       if $\mathop{\mathtt{value}}(CS_{level},M)=\mathop{\mathtt{true}}$ then
11             if $assignment[x_{level}]$ not in $\textnormal{{Solve}}(M)$ then
12                   $assignment[x_{level}]\leftarrow$ one element in $\textnormal{{Solve}}(M)$
13            $M\leftarrow[M;x_{level}\mapsto assignment[x_{level}]]$
14            $level++$
15             $maxlevel\leftarrow\max(level,maxlevel)$
16            if $level>n$ then return ‘SAT’
17             if GoTo2dLS? then
18                   $(status,\alpha,numJump)\leftarrow\allowbreak 2d\text{-LS}(F\mid_{x_{1}=assignment[x_{1}],\ldots,x_{level-1}=assignment[x_{level-1}]},MaxRestart_{2},MaxJump,m)$
19                  if $status=$‘SAT’ then return ‘SAT’
20                   else
21                         for i from $level$ to $n$ do $assignment[x_{i}]\leftarrow\alpha[x_{i}]$
22
23                        $numFailCells\leftarrow numFailCells+numJump$
24
25
26       else
27             if there exists a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{false}}$ then  $mcstatus\leftarrow(\text{`UNSAT'},c)$
28             else if there exists a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{undef}}$ and only one literal $\ell$ in clause $c$ such that $\mathop{\mathtt{value}}(\ell,M)=\mathop{\mathtt{undef}}$ then  $mcstatus\leftarrow(\text{`Propagate'},\ell,c)$
29             else
30                   choose a clause $c$ in $CS_{level}$ such that $\mathop{\mathtt{value}}(c,M)=\mathop{\mathtt{undef}}$
31                  choose a literal $\ell$ in $c$ such that $\mathop{\mathtt{value}}(\ell,M)=\mathop{\mathtt{undef}}$
32                  $mcstatus\leftarrow(\text{`Decide'},\ell,c)$
33
34            if $mcstatus[1]\neq$ ‘UNSAT’  then
35                   $\ell\leftarrow mcstatus$[2], $c\leftarrow mcstatus$[3]
36                  if $\mathop{\mathtt{Consistent}}(\ell,M)$  then
37                         if $status[1]=$ ‘Decide’ then  $M\leftarrow[M;\ell]$
38                         else if $status[1]=$ ‘Propagate’ then  $M\leftarrow[M;c\to\ell]$
39
40                   else
41                         $minCore\leftarrow$ minimal conflicting core of $M$ and $\ell$ on the level $level$
42                        $cell\leftarrow\mathop{\mathtt{explain}}(minCore,(assignment[x_{1}],\ldots,assignment[x_{level-1}]))$
43                        $lemma\leftarrow\neg\left(cell\land(\bigwedge_{\ell^{\prime}\in minCore}\ell^{\prime})\land\ell\right)$
44                        $learnClauses\leftarrow learnClauses\cup\{lemma\}$
45                        $CS_{level}\leftarrow CS_{level}\cup\{lemma\}$
46                        $M\leftarrow[M;lemma\to\neg\ell]$
47                        if $status[1]=$ ‘Propagate’ then
48                               $numFailCells\leftarrow numFailCells+1$
[…]
50
51
52
53            if $mcstatus[1]=$ ‘UNSAT’  then
54                   $c\leftarrow mcstatus[2]$, $lemma\leftarrow\mathop{\mathtt{resolve}}(c,M,level)$
55                  if $lemma$ is empty then return ‘UNSAT’
56
57                  $learnClauses\leftarrow learnClauses\cup\{lemma\}$
58                  if $x_{level}$ appears in $lemma$ then
59                         $CS_{level}\leftarrow CS_{level}\cup\{lemma\}$
60                        $\ell^{*}\leftarrow$ the last decided literal in $M$ satisfying $\ell^{*}\in\mathop{\mathtt{atoms}}(lemma)$
61                        delete the decided literal $\ell^{*}$ and all subsequent terms from $M$
62
63                  else
64                         $tmpLevel\leftarrow$ maximal $i$ of $x_{i}$ appearing in $lemma$
65                        $CS_{tmpLevel}\leftarrow CS_{tmpLevel}\cup\{lemma\}$
66                        delete the assignment $x_{tmpLevel}\mapsto assignment[x_{tmpLevel}]$ and all subsequent terms from $M$
67                        $level\leftarrow tmpLevel$
68
69
70       if $numFailCells>max\_numFailCells$ then
71            return OpenCAD$(F,learnClauses)$
72
Algorithm 2 The Hybrid Framework
Inspired by the work of Zhang et al. (Zhang et al., 2024) on combining CDCL(T) and local search for SMT-NIA, we propose a hybrid framework for SMT-NRA that integrates $2d$-LS (see Alg. 1), MCSAT (de Moura and Jovanović, 2013) and OpenCAD (Han et al., 2014) in this section. Sect. 4.1 presents an overview of the hybrid framework, while Sect. 4.2 provides a detailed explanation of our MCSAT implementation.
4.1. Overview of the Hybrid Framework
The hybrid framework (also see Alg. 2) consists of the following three main stages.
Stage 1: $\boldsymbol{2d}$-LS. Given a polynomial formula such that every relational operator appearing in it is ‘$<$’ or ‘$>$’, the first stage (Alg. 2, line 2–line 2) tries to solve the satisfiability by calling the $2d$-LS algorithm, that is Alg. 1. The reason for employing the local search algorithm first lies in its lightweight and incomplete property. If it successfully finds a model, the solving time is typically short. If it fails, the subsequent stages apply complete solving algorithms. In addition, two key outputs from $2d$-LS are maintained for later stages: the final cell-jump location and the number of cell-jump steps (Alg. 2, line 2). The final cell-jump location provides candidate variable assignments for the second stage, while the number of cell-jump steps helps to estimate the number of unsatisfiable cells for the input formula, which will be used in the third stage.
Stage 2: $\boldsymbol{2d}$-LS-Driven MCSAT. The second stage (Alg. 2, line 2–line 2) adopts an MCSAT framework as its foundational architecture. Recall that original MCSAT framework (de Moura and Jovanović, 2013) assigns variables one-by-one without imposing constraints on variable assignments. For example, for the theory of nonlinear real arithmetic, every variable can be assigned an arbitrary rational number. However, in Alg. 2, following each variable assignment in the MCSAT framework, we add a heuristic condition (Alg. 2, line 2) to determine whether the input formula may be satisfiable under the current variable assignments. If the formula is determined to have a high probability of being satisfiable, we invoke the $2d$-LS algorithm (i.e., Alg. 1) for the formula with all variables assigned so far replaced with their assigned values in MCSAT.
  1. (1)
    If the output is ‘SAT’, combining current variable assignments in MCSAT and the model found by Alg. 1, we obtain a model for the original input formula.
  2. (2)
    Otherwise, we use the final cell-jump location of Alg. 1 as candidate variable assignments for the unassigned variables in MCSAT. (This choice is motivated by the fact that $2d$-LS has performed multiple local modifications near this assignment without success, suggesting its proximity to critical constraints that may lead to conflicts. From a geometric perspective, there may be many unsatisfiable CAD cells near the CAD cell where this assignment is located. Therefore, the MCSAT can learn a lemma covering multiple adjacent CAD cells from this assignment, thereby improving the efficiency of MCSAT.)
This stage employs a local search algorithm to efficiently solving sub-formula satisfiability under partial assignments and drive variable assignments in the MCSAT framework, achieving an organic integration of the two methods. Hence, we call the stage “$2d$-LS-driven MCSAT”. For further implementation details of our MCSAT framework, please refer to Sect. 4.2. Additionally, to prepare for the third stage, we collect learned clauses generated during the MCSAT procedure (see Alg. 2, line 2 and line 2), and updates the estimation for the number of unsatisfiable cells (see Alg. 2, line 2).
Stage 3: OpenCAD. Recall that in the first two stages, we estimate the number of unsatisfiable cells for the input formula (note this represents a rough approximation, rather than the exact number of unsatisfiable cells). During the second stage, once the estimation exceeds a predetermined threshold (see Alg. 2, line 2), the algorithm transitions to this stage by invoking the OpenCAD procedure. OpenCAD is a complete method for deciding satisfiability of quantifier-free formulas comprising exclusively strict inequality constraints. The integration of OpenCAD into our hybrid framework is motivated by the complementary strengths of the two complete methods: MCSAT is a CDCL-style search framework, demonstrating superior performance on determining the satisfiability of satisfiable formulas or unsatisfiable ones dominated by Boolean conflicts, while OpenCAD specializes in efficiently handling unsatisfiable formulas dominated by algebraic conflicts (refer to Example 4.1 for a more detailed explanation). Moreover, we provide learned clauses from the MCSAT procedure for the OpenCAD invocation. The lifting process of OpenCAD terminates immediately upon detecting any low-dimensional region that violates either the input formula or these learned clauses.
Example 4.1.
Consider the polynomial formula $F>0$, where the polynomial
$F=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2})^{2}-4(x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2}+x_{3}^{2}x_{4}^{2}+x_{4}^{2}x_{5}^{2}+x_{5}^{2}x_{1}^{2}).$
The Boolean structure of this formula is remarkably simple, with a single polynomial constraint. However, the algebraic structure of the only polynomial in the formula is highly complex. Given the polynomial in $F$ as input, a CAD algorithm partitions $\mathbb{R}^{5}$ into approximately $2000$ cells, on which the polynomial has constant sign, either $+$, -- or 00. In the MCSAT framework, it is tedious to encode every unsatisfiable cell of $F$. Thus, for satisfiability solving of formula $F$, OpenCAD demonstrates superior efficiency compared to MCSAT.
4.2. MCSAT Implementation
The original MCSAT framework is formulated by transition relations between search states, as described in (Jovanović and De Moura, 2013; de Moura and Jovanović, 2013). In order to present our hybrid framework more clearly, we provide a pseudocode representation in Alg. 2, where the black lines correspond to our implementation of the original MCSAT procedure, and the blue lines represent the extensions introduced by the hybrid framework.
Let $F$ be a polynomial formula, $c$ be a clause, $\ell$ be an atom, and $f$ be a polynomial in $\mathbb{Q}[x_{1},\ldots,x_{n}]$. We denote by $\mathop{\mathtt{clauses}}(F)$, $\mathop{\mathtt{atoms}}(c)$, $\mathop{\mathtt{poly}}(\ell)$, and $\mathop{\mathtt{vars}}(f)$ the set of all clauses in formula $F$, the set of all atoms in clause $c$, the polynomial appearing in atom $\ell$, and the set of all variables appearing in polynomial $f$, respectively.
The level of variable $x_{i}$, polynomial $f$, atom $\ell$, and clause $c$ is defined as follows: $\mathop{\mathtt{level}}(x_{i})=i$, $\mathop{\mathtt{level}}(f)=\max(\{\mathop{\mathtt{level}}(x_{i})\mid x_{i}\in\mathop{\mathtt{vars}}(f)\})$, $\mathop{\mathtt{level}}(\ell)=\mathop{\mathtt{level}}(\mathop{\mathtt{poly}}(\ell))$, and $\mathop{\mathtt{level}}(c)=\max(\{\mathop{\mathtt{level}}(\ell)\mid\ell\in\mathop{\mathtt{atoms}}(c)\})$.
Suppose the input polynomial formula of Alg. 2 is $F$. We explain the notations used in Alg. 2.
  • $\bullet$
    For $i=1,\ldots,n$, $CS_{i}$ (see Alg. 2, line 2) represents the set of clauses in $F$ at level $i$.
  • $\bullet$
    The trail $M$ (see Alg. 2, lines 2, 2, 22 and 2) is a sequence of decided literals, propagated literals and variable assignments. A decided literal, denoted by $\ell$, meaning that $\ell$ is assumed to be true. A propagated literal, denoted by $c\rightarrow\ell$, meaning that $\ell$ is implied to be true in the current trail by clause $c$. The trail $M$ takes the following level-increasing form
    $[M_{1};x_{1}\mapsto assignment[x_{1}];M_{2};x_{2}\mapsto assignment[x_{2}];\ldots;x_{i}\mapsto assignment[x_{i}];M_{i}],$
    where for $k=1,\ldots,i$, the sequence $M_{k}$ does not contain any variable assignments, and every element in $M_{k}$ is at level $k$, i.e.$\mathop{\mathtt{level}}(\ell)=k$, whether it is a decided literal $\ell$ or a propagated literal $c\rightarrow\ell$.
  • $\bullet$
    $\mathop{\mathtt{value}}(CS_{i},M)$ is the value (true/false/undef) of the conjunction of clauses in $CS_{i}$ by replacing literals in $M$ to be true, and negation of literals in $M$ to be false. Similarly, $\mathop{\mathtt{value}}(c,M)$ is the value of the clause $c$, and $\mathop{\mathtt{value}}(l,M)$ is the value of the literal $l$.
  • $\bullet$
    $\texttt{Solve}(M)$ is the solution interval of $x_{level}$ restricting literals in $M$ to be true, where $M$ has $level$ levels. (Note that assignments of $x_{1},\ldots,x_{level-1}$ are fixed by $M$, so the solution interval of $x_{level}$ can be derived).
  • $\bullet$
    $\texttt{Consistent}(\ell,M)$ is the consistency (true/false) of the literal $\ell$ and literals in $M$, i.e., whether $x_{level}$ has a solution under the constraints that $\ell$ is true and literals in $M$ are true.
  • $\bullet$
    The “minimal conflicting core of $M$ and $\ell$ at the level $level$” refers to a minimal subset of literals at level $level$ in $M$ which is inconsistent with $\ell$.
  • $\bullet$
    The function $\texttt{explain}(minCore,a:=(assignment[x_{1}],\ldots,assignment[x_{level-1}]))$ constructs a cell (by the single-cell projection operator $\mathop{\mathtt{Proj}}$, see Definition 2.4) of the set of polynomials in $minCore$, denoted by $P\subseteq\mathbb{Q}[x_{1},\ldots,x_{level}]$, and the sample point $a\in\mathbb{R}^{level-1}$. Formally, $\texttt{explain}(minCore,a)=\mathop{\mathtt{Proj}}(P,x_{level},a)$.
  • $\bullet$
    $\texttt{resolve}(c,M,level)$ is the resolution of the clause $c$ and the conjunction of literals in $M$ at the level $level$.
Next, we explain the our implementation of MCSAT. MCSAT combines model constructing with conflict driven clause learning. The core idea of MCSAT is levelwise constructing a model by assigning variables according to a fixed order, and generating lemmas when encountering conflicts to avoid redundant search. The input of MCSAT is a polynomial formula $F$ with variables $x_{1},\ldots,x_{n}$, such that the relational operator of every atom is ‘$<$’ or ‘$>$’. The output of MCSAT is ‘SAT’ if $F$ is satisfiable, or ‘UNSAT’ if $F$ is unsatisfiable. MCSAT has the following four core strategies.
Assign (line 2–line 2): If clauses in $CS_{level}$ are evaluated to be true, then the variable $x_{level}$ is assigned to be one element in the interval $\texttt{Solve}(M)$, and MCSAT’s search enters the next level. If all variables are assigned, MCSAT returns ‘SAT’.
Status Update (line 2–line 2, line 2–line 2): The status of MCSAT $mcstatus$ is updated after deciding the value of an undefined literal in a clause (‘Decide’), propagating an undefined literal in a clause (‘Propagate’), or determining $F$ is unsatisfiable if the value of a clause is false (‘UNSAT’).
Consistency Check (black lines in line 2–line 2): The decided or propagated literal $\ell$ is checked if it is consistent with $M$. If so, the literal is appended to $M$. Otherwise, MCSAT learns a lemma $lemma$ which is the negation of the conjunction of three components. The first component $cell$ is a cell constructed by the single-cell projection operator, where polynomials involved in $minCore$ are sign-invariant in this cell; the second component $\land_{\ell^{\prime}\in minCore}\ell^{\prime}$ consists of literals in $minCore$; the third component is the literal $\ell$. When the second and the third components are satisfied, every assignment in $cell$ is unsatisfied due to the inconsistency of $\ell$ with $M$. Therefore, the lemma $lemma$ represents the three components cannot hold simultaneously is learnt. Then, $lemma$ is added to $CS_{level}$, and $lemma\rightarrow\neg\ell$ is appended to $M$. If the literal $\ell$’s inconsistency with $M$ stems from propagation, MCSAT updates its status $mcstatus$ to ‘UNSAT’.
Conflict Resolve (black lines in line 2–line 2): When the status of MCSAT is ‘UNSAT’, MCSAT learns a lemma $lemma$ by resolution. If $lemma$ is empty, MCSAT returns ‘UNSAT’. MCSAT backtracks by undoing the last decided literal. If $x_{level}$ appears in $lemma$, MCSAT backtracks to the last decided literal $\ell^{*}$ in $M$ that is an atom of $lemma$; otherwise, MCSAT backtracks to the last variable assignment in $lemma$.
Overall, the MCSAT algorithm operates through four core strategies working in coordination. Beginning with Assign that assigns the variable when all clauses at the current level evaluate to true, the algorithm may terminate with ‘SAT’ if all variables are successfully assigned. When assignment fails, Status Update yields either ‘Decide’, ‘Propagate’, or ‘UNSAT’ statuses. For ‘Decide’ or ‘Propagate’ statuses, Consistency Check verifies the consistency of the decided or propagated literal with the trail $M$, and any detected inconsistency triggers the lemma learning that involves the single-cell projection operator ($\mathop{\mathtt{Proj}}$, see Definition 2.4). Besides, inconsistency under the status ‘Propagate’ leads to the status ‘UNSAT’ by updating in Status Update. For ‘UNSAT’ statuses, Conflict Resolve conducts lemma learning (an empty lemma makes the algorithm terminate with ‘UNSAT’), and then executes non-chronological backtracking driven by the lemma. This process repeats iteratively until termination.
5. Experiments
5.1. Experiments Setups
5.1.1. Environment
In this paper, all experiments are conducted on 8-Core 11th Gen Intel(R) Core(TM) i7-11700@2.50GHz with 16GB of memory under the operating system Ubuntu 24.04 (64-bit). Each solver is given a single attempt to solve each instance, with a time limit of 1200 seconds.
5.1.2. Benchmarks
  • $\bullet$
    SMTLIB: This refers to a filtered subset of the QF_NRA benchmark from the SMT-LIB [^1] http://smtlib.cs.uiowa.edu/, [^2] https://zenodo.org/communities/smt-lib/ standard benchmark library, containing all instances with only strict inequalities. This selection matches the input requirements of our hybrid framework (see Alg. 2). There are $2050$ instances in this benchmark.
  • $\bullet$
    Rand Inst: This refers to random instances, which are generated by the random polynomial formula generating function $\mathbf{rf}$ and parameters in (Li et al., 2023, Sect. 7.2), i.e.$\mathbf{rf}$({30,40}, {60,80},{20,30},{10,20},{20,30},{40,60},{3,5}). These instances are almost always satisfiable. Unlike SMTLIB’s abundance of unit clauses and linear constraints, the random instances are more complicated in the Boolean structure (i.e., every clause has at least 3 atoms) and highly nonlinear (i.e., the degree of every polynomial is at least 20). There are 100 instances in this benchmark.
  • $\bullet$
    Spec Inst: This refers to specific instances, which are all instances in (Li and Xia, 2020, Sect. 5), including Han_$\_n$, P, Hong_$\_n$, Hong2_$\_n$ and C_$\_n\_r$. These instances have particular mathematical properties that pose challenges for conventional solvers. There are 21 instances in this benchmark.
5.1.3. Implementation
[…]
5.2. Competitors
  • $\bullet$
    Our Main Solver: HELMS is the hybrid solver that integrates $2d$LS, MCSAT and OpenCAD according to Alg. 2.
  • $\bullet$
    Base Solvers: HELMS integrates two base solvers. The first base solver LS is a local search solver using the original cell-jump operation (configuration from (Li et al., 2023)). The solver name coincides with that of the original local search algorithm it implements. For clarity, ‘LS’ hereafter denotes the original local search solver. The second base solver LiMbS is an MCSAT solver following the description in Sect. 4.2. (As LS only accepts inputs with strict inequalities, LiMbS and HELMS inherit this input restriction.)
  • $\bullet$
    SOTA Solvers: Four SOTA solvers from recent SMT Competitions (SMT-COMP) are Z3 (version 4.13.4), CVC5 (version 1.2.0), MathSAT5 (version 5.6.11) and Yices2 (version 2.6.5).
5.2.1. Indicators
[…]
5.3. Performance Comparison between HELMS and SOTA Solvers
Table 1. The number of instances solved by HELMS, base solvers and SOTA solvers on three benchmarks.
Benchmark #INST HELMS LS LiMbS Z3 CVC5 MathSAT5 Yices2
SMTLIB #SAT 1503 1503 1472 1502 1503 1482 812 1496
#UNSAT 547 542 - 538 536 535 315 538
#ALL 2050 2045 - 2040 2039 2017 1127 2034
Rand Inst #SAT 100 100 100 71 16 0 1 23
Spec Inst #SAT 7 7 4 7 7 4 1 3
#UNSAT 14 14 - 14 5 4 2 3
#ALL 21 21 - 21 12 8 3 6
Total #SAT 1610 1610 1576 1580 1526 1486 814 1522
#UNSAT 561 556 - 552 541 539 317 541
#ALL 2171 2166 - 2132 2067 2025 1131 2063
(a) Speed comparison between HELMS and Z3 on three benchmarks.
(b) Speed comparison between HELMS and CVC5 on three benchmarks.
(c) Speed comparison between HELMS and MathSAT5 on three benchmarks.
(d) Speed comparison between HELMS and Yices2 on three benchmarks.
Figure 4. Speed comparison between HELMS and SOTA solvers on three benchmarks - SMTLIB (red), Rand Inst (blue), and Spec Inst (green). For the heatmap corresponds to each SOTA solver, the vertical axis represents the runtime intervals of the SOTA solver, while the horizontal axis indicates the ratio of the SOTA solver’s runtime to that of HELMS — reflecting how many times faster HELMS is compared to the SOTA solver. The first column (‘Total’) shows the total number of instances solved by the SOTA solver within each runtime interval. The remaining columns (‘1+’, ‘10+’, ‘100+’, and ‘1000+’) indicate the number of instances where HELMS is at least 1, 10, 100, or 1000 times faster than the SOTA solver, respectively.
For benchmark SMTLIB:
  • $\bullet$
    In terms of the number of solved instances, as shown in Tab. 1, row ‘SMTLIB’, HELMS solves 2045 out of 2050 instances, outperforming SOTA solvers. While both HELMS and Z3 solve the most SAT instances, HELMS additionally solves the most UNSAT instances.
  • $\bullet$
    In terms of runtime, as shown in Fig. 4-4, columns ‘1+’ (red), HELMS is faster than each SOTA solver on approximately 1000 instances. While the runtime of SOTA solvers cluster in intervals 0.01-0.1s and 0.1-1s, HELMS achieves 10 times speedup for many instances in these intervals, indicating HELMS has speed advantage on common instances in SMTLIB. Instances correspond to rows ‘1000-1200’ in Fig. 4-4 are typically instances that SOTA solvers timeout. HELMS achieves 100 or 1000 times speedup on many such instances, demonstrating HELMS’s ability to solve computationally demanding problems.
For benchmark Rand Inst:
  • $\bullet$
    In terms of the number of solved instances, as shown in Tab. 1, row ‘Rand Inst’, HELMS solves all 100 instances, demonstrating superior performance compared to SOTA solvers. Among the SOTA solvers, Yices2 performs the best yet solves only 23 instances, while CVC5 fails to solve any. These results highlight HELMS’s distinct advantage in handling SAT problems that combine complex Boolean structure with highly nonlinear constraints. While SOTA solvers typically excel at problems with intricate Boolean structure, their performance significantly degrades when confronted with highly nonlinear constraints. The superior performance of HELMS stems from the cell-jump operation in LS, which effectively identifies and jumps to assignments that flip the signs of high-degree polynomials. This advantageous property is preserved in the $2d$-cell-jump operation in the $2d$-LS employed by HELMS, maintaining its effectiveness for handling high-degree algebraic constraints.
  • $\bullet$
    In terms of runtime, as shown in Fig. 4-4, rows ‘100-1000’ and ‘1000-1200’ (blue), SOTA solvers spend 100-1000s and 1000-1200s for major random instances. As shown in Fig. 4-4, columns ‘1000+’ (blue), HELMS is generally 1000 times faster than SOTA solvers, representing that HELMS is efficient in this benchmark. HELMS’s efficiency for highly nonlinear problems is because $2d$-cell-jump in $2d$-LS may ‘jump’ to a solution faster for a formula with high-degree polynomials.
For benchmark Spec Inst:
  • $\bullet$
    In terms of the number of solved instances, as shown in Tab. 1, row ‘Spec Inst’, HELMS solves all 21 instances. In contrast, SOTA solvers fail to complete many instances within the time limit, regardless of their satisfiability status (SAT or UNSAT). HELMS excels on instances with mathematical properties that challenge SOTA solvers. The explanation is as follows. For SAT instances, $2d$-cell-jump in $2d$-LS is effective in more complicated algebraic problems; For UNSAT instances, (1) the sample-cell projection operator in MCSAT generates larger unsatisfiable cells to be excluded, and (2) OpenCAD is effective in handling algebraic conflicts.
  • $\bullet$
    In terms of runtime, as shown in Fig. 4-4, rows ‘1000-1200’ (green), SOTA solvers spend 1000-1200s for many specific instances. As shown in Fig. 4-4, columns ‘10+’ and ‘100+’ (green), HELMS is generally 10 or 100 times faster than SOTA solvers. This performance gap indicates that while these problems remain computationally difficult, HELMS can solve them within practical time limits.
For all benchmarks:
  • $\bullet$
    In terms of the number of solved instances, as shown in Tab. 1, row ‘Total’, HELMS solves most SAT and UNSAT instances, outperforming SOTA solvers, especially MathSAT5.
  • $\bullet$
    In terms of runtime, as shown in Fig. 5, HELMS is competitive with SOTA solvers. The advantage of HELMS is particularly evident on SAT instances. The bias towards SAT instances is due to the use of $2d$-LS in Stage 1 (see Sect. 4), which effectively accelerates model search. The cumulative distribution function (CDF) analysis in Fig. 6 reveals HELMS’s overall runtime advantage, with its curve positioned mainly to the left of those for Z3, CVC5, and Yices2. The section of the curve to the left of those for Z3, CVC5, and Yices2 indicates that HELMS is slower than them on UNSAT cases. A comparison between Fig. 6 and Fig. 6 shows that while HELMS’s performance on SAT instances follows a trend similar to SOTA solvers, its CDF curve for all instances exhibits a depression in the upper left corner. This phenomenon occurs because Stage 1 processing delays the determination of UNSAT cases - UNSAT results can only be confirmed after Stage 1 completion, leading to right-shifted runtime measurements for these instances.
(a) Runtime comparison of HELMS and Z3.
(b) Runtime comparison of HELMS and CVC5.
(c) Runtime comparison of HELMS and MathSAT5.
(d) Runtime comparison of HELMS and Yices2.
Figure 5. Runtime comparison of HELMS and SOTA solvers in all benchmarks. Each data point in the plot represents one instance, with orange plus symbols (’+’) denoting SAT instances and blue minus symbols (’-’) denoting UNSAT instances. Data points above/below the diagonal represent instances where HELMS is faster/slower, respectively. Points positioned at the top/right boundary (1200s) indicate instances where the SOTA solver or HELMS timeouts, respectively.
(a) All instances.
(b) SAT instances.
Figure 6. CDFs of HELMS and SOTA solvers on all benchmarks. The position of the CDF curve to the left signifies a more rapid solver performance, while a higher curve indicates a greater number of instances that the solver is capable of solving.
5.4. Effectiveness of Proposed Strategies
5.4.1. Effectiveness of Extending Local Search with MCSAT
To show the effectiveness of extending local search with MCSAT (see Sect. 3), we compare the performance of $2d$-LS and LS on SAT instances in all benchmarks.
Figure 7. Runtime of $2d$-LS and LS on SAT instances in all benchmarks. Points above/below the diagonal denote instances where $2d$-LS/LS is faster.
  • $\bullet$
    In terms of the number of solved instances, as shown in Tab. 1, $2d$-LS solves all SAT instances, while LS fails to solve 3 instances in Spec Inst. This performance advantage confirms that by extending the one-dimensional search space (LS, cell-jump) to two dimensions ($2d$-LS, $2d$-cell-jump) enables access to a broader solution space, reaching models that are inaccessible in the one-dimensional case
  • $\bullet$
    In terms of runtime, as shown in Fig. 7, for instances that LS spends less than 0.1s, $2d$-LS solves nearly all such instances within 0.1s as well. However, for instances that LS spends more than 0.1s, $2d$-LS almost completely surpasses LS. These results validate that the $2d$-cell-jump operation enhances model-search efficiency compared to the cell-jump operation.
5.4.2. Effectiveness of the Hybrid Framework
To show the effectiveness of the hybrid framework (see Sect. 4), we evaluate HELMS in the following aspects.
  1. (1)
    HELMS solves more instances than its base solvers by integrating their respective strengths.
    • $\bullet$
      For benchmark SMTLIB, as shown in Tab. 1, row ‘SMTLIB’, LS solves 1472 SAT instances, HELMS solves more SAT instances than LS (contributed by $2d$-LS and MCSAT); LiMbS solves 1502 SAT instances, HELMS solves one more SAT instance than LiMbS (contributed by $2d$-LS); LiMbS solves 538 UNSAT instances, HELMS solves four more UNSAT instances than LiMbS (contributed by OpenCAD).
    • $\bullet$
      For benchmark Rand Inst, as shown in Tab. 1, row ‘Rand Inst’, LS solves all 100 instances, whereas LiMbS only solves 71 of them. HELMS also solves all of them. These types of instances are where LS excels, and HELMS effectively incorporates this strength.
    • $\bullet$
      For benchmark Spec Inst, as shown in Tab. 1, row ‘Spec Inst’, LiMbS solves all 21 instances, while LS solves only 4 SAT instances. HELMS solves all of them, benefiting from LiMbS’s capabilities. These types of instances are where LiMbS excels, and HELMS effectively incorporates this strength.
  2. (2)
    Each stage of HELMS contributes effectively to overall performance.
    Table 2. Stages that HELMS ends in solving instances on three benchmarks.
    Benchmark #INST HELMS Stage 1 Stage 2 Stage 3
    ($2d$-LS) ($2d$-LS-Driven MCSAT) (OpenCAD)
    SMTLIB #SAT 1503 1503 905 598 0
    #UNSAT 547 542 0 535 7
    #ALL 2050 2045 905 1133 7
    Rand Inst #SAT 100 100 99 1 0
    Spec Inst #SAT 7 7 7 0 0
    #UNSAT 14 14 0 11 3
    #ALL 21 21 7 11 3
    Total #SAT 1610 1610 1011 599 0
    #UNSAT 561 556 0 546 10
    #ALL 2171 2166 1011 1145 10
    • $\bullet$
      Stage 1 ($2d$-LS) is effective since it contributes to determining the satisfiability of 1011 instances, as shown in Tab. 2. This stage is important for SAT instances in all benchmarks. Most SAT instances in SMTLIB are determined by Stage 1, indicating HELMS leverages $2d$-LS’s strength in quickly finding models. Nearly all SAT instances in Rand Inst and Spec Inst are also solved at this stage, demonstrating HELMS leverages $2d$-LS’s strength in handling highly nonlinear polynomials and polynomials.
    • $\bullet$
      Stage 2 ($2d$-LS-Driven MCSAT) is effective since it contributes to determining the satisfiability of 1145 instances, as shown in Tab. 2. This stage is important for UNSAT instances dominated by Boolean conflicts. Most UNSAT instances in SMTLIB are determined by Stage 2, indicating HELMS leverages the strength of MCSAT to solve common UNSAT problems efficiently. Most UNSAT instances in Spec Inst are determined by Stage 2, demonstrating HELMS leverages the sample-cell projection operator’s effect in problems with difficult mathematical properties.
    • $\bullet$
      Stage 3 (OpenCAD) is effective since it contributes to determining the satisfiability of 10 instances, as shown in Tab. 2. This stage is important for UNSAT instances dominated by algebraic conflicts. A small portion of UNSAT instances in SMTLIB and Spec Inst are hard and detected to have complex algebraic structure. The detection is done by Stage 2, and this information hints HELMS to switch to Stage 3. HELMS leverages OpenCAD’s strength in algebraic reasoning.
  3. (3)
    Each stage of HELMS is necessary to overall performance.
    We conduct an ablation study by modifying HELMS to the following variants.
    • $\bullet$
      V1: Removing Stage 1.
    • $\bullet$
      V2: Removing Stage 2.
    • $\bullet$
      V3: Removing Stage 3.
    • $\bullet$
      V4: Removing Stage 1 and disabling $2d$-LS (Alg. 2, line 2–line 2) in Stage 2.
    • $\bullet$
      V5: Disabling using the final cell-jump location of Alg. 1 as candidate varible assignments for the unassigned variables in MCSAT (see Sect. 4.1, Stage 2, (2)).
    Table 3. The number of instances solved by HELMS and its variants on three benchmarks.
    Benchmark #INST HELMS V1 V2 V3 V4 V5
    SMTLIB #SAT 1503 1503 1503 1503 1503 1503 1503
    #UNSAT 547 542 542 449 536 541 541
    #ALL 2050 2045 2045 1952 2039 2044 2044
    Rand Inst #SAT 100 100 34 100 100 15 100
    Spec Inst #SAT 7 7 5 7 7 5 7
    #UNSAT 14 14 13 5 14 13 14
    #ALL 21 21 18 12 21 18 21
    Total #SAT 1610 1610 1542 1610 1610 1523 1610
    #UNSAT 561 556 555 454 550 554 555
    #ALL 2171 2166 2097 2160 2067 2077 2165
    In terms of the number of solved instances, as shown in Tab. 3, all of the variants are worse than HELMS in certain benchmarks. V1 solves 66 fewer random instances, and 2 fewer specific instances than HELMS. The solving of these instances in HELMS are all contributed by the $2d$-LS in Stage 2, indicating Stage 1 is necessary for highly nonlinear SAT problems. V2 solves 102 fewer UNSAT instances than HELMS, indicating Stage 2 is necessary for UNSAT instances. This also shows that the concatenation of $2d$-LS and OpenCAD cannot effectively solve UNSAT cases. V3 solves 6 fewer UNSAT instances than HELMS, indicating Stage 3 is necessary for UNSAT problems dominated by algebraic conflicts. V4 solves 19 fewer random instances than V1, signifying the necessity of $2d$-LS in Stage 2. Also, V4 misses one UNSAT instance in SMTLIB that HELMS solves via OpenCAD. The reason for this phenomenon is that when the $2d$-LS in Stage 2 is disabled, the number $numJump$ (see Alg. 2) increases much slower, such that the heuristic condition to switch to OpenCAD cannot be satisfied. It indicates that $2d$-LS in Stage 2 is necessary to detect problems with complex algebraic structure, i.e., the number of CAD cells is large. V5 solves one fewer UNSAT instance than HELMS, because of the initial variable assignments for MCSAT in V5 lack $2d$-LS guidance, proving the effectiveness of this strategy.
    (a) Runtime comparison of HELMS and V1.
    (b) Runtime comparison of HELMS and V2.
    (c) Runtime comparison of HELMS and V3.
    (d) Runtime comparison of HELMS and V4.
    (e) Runtime comparison of HELMS and V5.
    Figure 8. Runtime comparison of HELMS and its variants in all benchmarks.
    In terms of runtime, as shown in Fig. 8 and Tab. 4, HELMS is superior over its variants across most instances. HELMS is faster than V1 on 79% (1273/1610) SAT instances, benefiting from the efficiency of $2d$-LS in handling SAT cases. On the contrary, HELMS is slower than V1 on most UNSAT instances, as Stage 1 cannot prove UNSAT. HELMS is faster than V2 on 96% (540/561) UNSAT instance, leveraging MCSAT’s effectiveness for UNSAT problems, including both regular UNSAT instances in SMTLIB, and UNSAT instances with special mathematical properties in Spec Inst. HELMS and V3 exhibit comparable runtime performance, except when V3 times out on particularly challenging UNSAT instances where OpenCAD is invoked. As evidenced by the comparison between V4 and V1, the $2d$-LS effectively determines when to transition to OpenCAD without introducing significant runtime overhead. HELMS is faster than V5 on 76% (1656/2171) instances, validating the necessity of using final assignments of $2d$-LS as initial assignments for MCSAT.
    Table 4. The number of instances that HELMS is faster than its variants on all benchmarks.
    Benchmark #INST V1 V2 V3 V4 V5
    SMTLIB #SAT 1503 1185 1214 1209 1413 1302
    #UNSAT 547 125 529 467 163 347
    #ALL 2050 1310 1743 1676 1576 1649
    Rand Inst #SAT 100 81 1 81 99 1
    Spec Inst #SAT 7 6 3 0 6 1
    #UNSAT 14 3 11 8 3 4
    #ALL 21 9 14 8 9 5
    Total #SAT 1610 1273 1219 1290 1519 1305
    #UNSAT 561 128 540 475 166 351
    #ALL 2171 1401 1759 1765 1685 1656
  4. (4)
    Additional HELMS configurations.
    (a) Full view.
    (b) Partial View.
    Figure 9. CDFs of HELMS adopting Strategy 1 with $len1=3,4,7$ and Strategy 2 with $len2=3,4$ on all benchmarks. The left figure presents the full time span (1200s), whereas the right figure shows an enlarged portion of the left figure.
    • $\bullet$
      Complexity bounds on rational numbers. In $2d$-LS, the complexity of rational-number assignments is restricted to a predetermined bound to prevent uncontrolled growth during iterations. We evaluate the following two bounding strategies.
      Strategy 1: Digit-length bounds on numerators and denominators. Let $\texttt{dig}(x)$ denote the number of digits in the positive integer $x$. For a rational number ±$\pm\frac{N}{D}$ (where $N$ and $D$ are coprime positive integers), if $\max(\texttt{dig}(N),\texttt{dig}(D))$ is greater than a predetermined number $len1$, then the rightmost $\min(\texttt{dig}(N),\texttt{dig}(D))-2$ digits from both terms are truncated. For example, $\frac{1234}{12345}$ is simplified to $\frac{12}{123}$, and $\frac{12345}{1234}$ is simplified to $\frac{123}{12}$.
      Strategy 2: Decimal rounding. Rounding rational numbers to $len2$ decimal places, where $len2$ is a predetermined number.
      We compare configurations with $len1=3,4,7$ and $len2=3,4$ in Fig. 9. While all settings solve the same number of instances, their efficiency are different. The $len1=4$ configuration under Strategy 1 proves the most efficient and is adopted in HELMS.
    Figure 10. Runtime of HELMS with Strategy 3 and Strategy 4 on instances in all benchmarks. Points above/below the diagonal denote instances where HELMS with Strategy 3/Strategy 4 is faster.
    • $\bullet$
      Preprocessing. The preprocessing in $2d$-LS derives variable intervals by analyzing constraints from all univariate linear atoms. For example, the atom $02x+3>$ 0 updates $x$’s domain from (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) to $(-\frac{3}{2},+\infty)$. Subsequent atoms involving $x$ will further intersect this interval with their derived constraints. This preprocessing prevents $2d$-LS from exploring easily detectable unsatisfiable regions. For a variable interval whose interval width is below $10^{-5}$, the variable is considered as approximately fixed, and its assignment would not be modified in $2d$-LS. If an instance is SAT when some of its variables are assigned to fixed values, then the instance is SAT. However, if an instance is UNSAT when some of its variables are assigned to fixed values, the instance may not be UNSAT. There are 702 instances with no more than two unfixed variables after preprocessing, which are all from SMTLIB benchmark. Recall that $2d$-LS performs $2d$-cell-jump, which implies the $2d$-LS for input with no more than two variables is the same with MCSAT algorithm. We compare following two strategies for instances that has no more than two unfixed variables after preprocessing.
      Strategy 3: Directly calling the MCSAT solver for the original instance.
      Strategy 4: Calling the MCSAT solver for the new problem of replacing the fixed variables in the instance with fixed values in their variable intervals. If the MCSAT solver returns SAT, then the original instance is SAT. Otherwise, the solver V1 (removing Stage 1 from HELMS) is invoked to solve the original problem.
      Fig. 10 demonstrates that both strategies solve the same number of instances, but Strategy 3 proves more efficient. A possible explanation is that the preprocessing, while tailored for $2d$-LS, offers limited utility to MCSAT. This is because MCSAT inherently handles univariate linear constraints efficiently, rendering the comparative advantage of Strategy 4 over Strategy 3 limited for SAT instances. For UNSAT cases, however, Strategy 4 incurs redundant computational cost, i.e., the MCSAT’s processing with the new problem is redundant. Consequently, Strategy 3 demonstrates superior overall efficiency, making it the preferred choice for HELMS.
6. Conclusion
In this paper, we have developed a hybrid SMT-NRA solver that integrates the complementary strengths of $2d$-LS, MCSAT and OpenCAD. First, we proposed a two-dimensional cell-jump operation, termed $2d$-cell-jump, which generalizes the key cell-jump operation in existing local search methods for SMT-NRA. Building on this, we proposed an extended local search framework, named $2d$-LS, which integrates MCSAT to realize $2d$-cell-jump in local search. To further improve MCSAT, we implemented the solver LiMbS that utilizes a recently proposed technique called the sample-cell projection operator, which is well suited for CDCL-style search in the real domain and helps guide the search away from conflicting states. Finally, we presented a hybrid framework that exploits the complementary strengths of MCSAT, $2d$-LS and OpenCAD. The hybrid framework consists of three stages. The first stage invokes $2d$-LS in an attempt to quickly find a satisfying assignment. The second stage adopts an MCSAT framework as a foundational architecture and triggers $2d$-LS when appropriate to suggest variable assignments for MCSAT, and signal the need to transition to the third stage. The third stage utilizes OpenCAD to handle unsatisfiable formulas that are dominated by algebraic conflicts. We implement our hybrid framework as the solver HELMS. Our experiments validate two key findings. First, HELMS is competitive with SOTA solvers in overall performance, and shows superior capability on problems with highly nonlinear constraints, complex Boolean structures, or particular mathematical properties. Second, the results verify the effectiveness of our proposed strategies, including the extension of local search with MCSAT and the design of the hybrid framework.